### v1

# Quantum Polymorphisms and Commutativity Gadgets

Abstract.
We introduce the concept of quantum polymorphisms to the complexity theory of non-local games. We use this notion to give a full characterisation of the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions. Prior to our work, a classification was only known in the Boolean case [Culf–Mastel, STOC’25]. As an application of our framework, we prove that the entangled CSP parameterised by odd cycles is undecidable. Furthermore, we establish a quantum version of Galois connection for entangled CSPs in the case of non-oracular quantum homomorphisms.
1. Introduction
In 1971, Cook’s introduction of $\mathsf{NP}$-completeness [cook1971complexity] provided a way to mathematically assess the inherent hardness of computational problems. Boolean satisfiability, graph colouring, linear equations over finite fields, and many other computatational problems of a seemingly distinct nature systematically began to fall into either the polynomial-time solvable or the $\mathsf{NP}$-complete buckets [Karp72], while intermediate complexity turned out to be a rarer phenomenon [ladner1975structure]. This fact soon sparked the hypothesis that the broad, natural class of computational problems involving satisfaction of constraints over finite alphabets—collectively known as (finite-domain) CSPs and encompassing the aforementioned problems and many others—could never exhibit the unnatural intermediate behaviour. $\operatorname{P}$ vs. $\mathsf{NP}$-complete’ dichotomies for fragments of CSPs started to emerge, primary examples being Schaefer’s [Schaefer78:stoc] and Hell–Nešetřil’s [HellN90] dichotomies for Boolean and graph CSPs, respectively. However, a unified framework for approaching the CSP Dichotomy Conjecture [Feder98:monotone] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [Jeavons97:closure, Jeavons98:algebraic]. By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [Bulatov17:focs] and, independently, Zhuk [Zhuk17_FOCS, Zhuk20:jacm] in 2017.
Seven years before the publication of Cook’s result, the work of Bell [bell1964einstein] originating from the Einstein–Podolsky–Rosen paradox [einstein1935can] demonstrated that quantum systems can display correlations unattainable within classical physics, thus ruling out local hidden-variable theories as realistic explanations for quantum mechanics. In its mathematical essence, Bell’s result can be formulated as a property of CSPs: There exist specific CSP languages for which the classical value and the entangled value can differ, where the latter is defined as the optimal success probability in a multiplayer interactive proof of the satisfiability of the CSP instance where the players have access to quantum resources. A primary question arising from Bell’s theorem concerns the computational complexity of determining the entangled value of a CSP. In particular, it is known that determining whether an instance of SAT is perfectly quantum satisfiable is undecidable (in fact, even the gap version of the problem is known to be undecidable, as a consequence of the celebrated $\operatorname{MIP^{*}=RE}$ theorem [ji2021mip]). On the other hand, examples of tractable CSP languages for which quantum and classical satisfiability coincide, which implies polynomial-time solvability of the corresponding entangled CSPs, are known [atserias_generalized_2019, ciardo_quantum_minion, bulatov2025satisfiability].
Nevertheless, unlike in the classical case, the complexity landscape of entangled CSPs remains largely unexplored. A natural strategy to shed some light on it is to transfer the algebraic machinery that allowed proving the (classical) CSP Dichotomy to the quantum setting. In particular, for classical CSPs, the concept of polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level. The root of this phenomenon is a Galois connection between relational clones (expressing the CSP language) and operation clones (expressing the family of polymorphisms). Via this connection, interreducibility between CSPs associated with different languages can be explained via inclusion (or weaker forms of the latter, in particular, so called height-1 clone homomorphisms) between their polymorphism clones.
To what extent can this machinery by transferred to the setting of entangled CSPs, and thus provide some understanding of their complexity and interreducibility? This question was first addressed at this level of generality by Ji in [ji2013binary]. In particular, that work identified a substantial obstacle against the prospect of quantising all types of algebraic CSP reductions known as gadgets. Ultimately, the obstacle is caused by the non-classical phenomenon of contextuality: Replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability. In the same paper, Ji provided a method to solve this issue for specific CSP languages, via the new notion of commutativity gadgets. On a high level, these are specific types of gadgets having no classical counterpart, that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction. For example, by showing that a specific triangular prism construction yields a commutativity gadgets for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding entangled CSPs, thus proving undecidability of quantum 3-colouring.
Later works generalised this idea to more general types of CSPs. In particular, commutativity gadgets for certain Boolean CSPs were constructed in [atserias_generalized_2019]. In a subsequent recent breakthrough [culf2024re], Culf and Mastel gave a complete characterisation of Boolean CSPs admitting commutativity gadgets. In the same work, the authors identified a sufficient property of CSP languages—which they called two-variable (non-)falsifiability—guaranteeing the existence of commutativity gadgets for the corresponding CSPs, and thus allows quantising classical reductions. Very recently, Culf, de Bruyn, Vernooij, and Zeman [Zeman] gave a different, necessary condition for the existence of commutativity gadgets, by showing that CSPs with non-classical quantum endomorphisms do not admit commutativity gadgets.
Despite this recent progress, a full characterisation of CSPs exhibiting commutativity gadgets is not available. In particular, [Zeman] identified odd cycles as a challenging open case for which the existence or non-existence of commutativity gadgets cannot be established via available methods.
Contributions
We give a complete characterisation of CSPs exhibiting commutativity gadgets. To that end, we introduce the notion of quantum polymorphisms, which lift classical polymorphisms to the setting of entangled CSPs. We show that quantum polymorphisms completely capture the existence of commutativity gadgets, by proving that a commutativity gadget exists if, and only if, the minion of quantum polymorphisms coincides with the quantum closure of the clone of standard polymorphisms. Through this characterisation, we prove that all odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable. This answers the question raised in [Zeman]. In addition, we use quantum polymorphisms to build commutativity gadgets for Boolean structures in a unified manner, thus giving a different proof of the result of [culf_re-completeness_2025]. Our characterisation works for both oracular and non-oracular CSPs. In the non-oracular setting, we additionally establish a quantum version of Galois connection that parallels the corresponding classical connection.
Structure of the paper
In Section 2, we give formal definitions of classical and entangled CSPs, and we list some basic results about quantum homomorphisms. Section 3 presents the main definition of this work—quantum polymorphims—as well as their use to characterise the existence of oracular commutativity gadgets for entangled CSPs. In Section 4 we show that the quantum polymorphisms of cliques are non-contextual, which gives a different proof of the undecidability of the corresponding entangled CSPs due to [ji2013binary, Zeman]. In section 5, we prove a general result on the consequence of contextuality in quantum homomorphisms, by showing that it causes the emergence of a specific non-orthogonality pattern in the underlying structures. We make use of this pattern in section 6, to prove that odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable. In Section 7, we address the non-oracular case, by showing that a similar characterisation to the one in Section 3 captures commutativity gadgets in the non-oracular setting. In addition, we establish therein a quantum version of Galois connection for the non-oracular quantum analogue of pp-definitions. Finally, in Section 8, we investigate the existence of commutativity gadgets for structures with a 2-element domain. We obtain that structures that do not admit the ternary majority as a polymorphism must have a commutativity gadget, and that apart from trivial cases, structures that do have the majority operation as a polymorphism do not have a commutativity gadget.
2. Preliminaries
2.1. Relational structures
For $\ell\in\mathbb{N}$, $\ell$] is the set $\{1,2,\dots,\ell\}$. A signature $\sigma$ is a finite set of relation symbols $R$, each with its arity $\operatorname{ar}(R)\in\mathbb{N}$. A relational structure $\mathbb{X}$ with signature $\sigma$ (in short, a $\sigma$-structure or simply a structure when $\sigma$ is implicit) consists of a finite set $X$ and a relation $R^{\mathbb{X}}\subseteq X^{\operatorname{ar}(R)}$ for each symbol $R\in\sigma$. All relational structures in this work shall be implicitly assumed to be finite. A (classical) homomorphism $f:\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ the the entries of $\mathbf{x}$). We denote the existence of a classical homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ by the notation $\mathbb{X}\to\mathbb{Y}$. Given a $\sigma$-structure $\mathbb{A}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq A^{r}$, we define the structure $\mathbb{A}_{S}$ as follows. The signature of $\mathbb{A}_{S}$ contains a unique $r$-ary symbol $R$ (which does not belong to $\sigma$), and the interpretation of $R$ in $\mathbb{A}_{S}$ is $R^{\mathbb{A}_{S}}=S$. For typographical reasons, we shall sometimes denote $\mathbb{A}_{S}$ by $(A;S)$. For a digraph $\mathbb{X}$, we let $\mathbb{X}^{u}$ be the undirected graph on the same vertex set of $\mathbb{X}$, having an undirected edge $\{x,y\}$ whenever $(x,y)$ or $(y,x)$ form an edge in $\mathbb{X}$. We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless directed graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing a pair $(x,y)$ if, and only if, $x$ and $y$ appear together in some tuple in some relation of $\mathbb{X}$. Given a $\sigma$-structure $\mathbb{Y}$, the constraint satisfaction problem parameterised by $\mathbb{Y}$ (in short, $\operatorname{CSP}(\mathbb{Y})$) is the following computational problem: [^1] This is the decision version of $\operatorname{CSP}(\mathbb{Y})$. In the search version, the problem is to find an explicit homomorphism $f:\mathbb{X}\to\mathbb{Y}$ for any input $\mathbb{X}$ such that $\mathbb{X}\to\mathbb{Y}$. The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [bulatov2005classifying]. Given as input a $\sigma$-structure $\mathbb{X}$, output $\mathrm{YES}$ if $\mathbb{X}\to\mathbb{Y}$, and $\mathrm{NO}$ if $\mathbb{X}\not\to\mathbb{Y}$. For $\ell\in\mathbb{N}$, we denote by $\mathbb{Y}^{\ell}$ the $\ell$-th direct power of $\mathbb{Y}$; i.e., $\mathbb{Y}^{\ell}$ is the $\sigma$-structure whose domain is $Y^{\ell}$ and whose relations are defined as follows: Given $R\in\sigma$ and any $\ell\times\operatorname{ar}(R)$ matrix $M$ whose rows are tuples in $R^{\mathbb{Y}}$, the columns of $M$ form a tuple in $R^{\mathbb{Y}^{\ell}}$. A polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$. By $\operatorname{Pol}(\mathbb{Y})$ (resp. $\operatorname{Pol}^{(\ell)}(\mathbb{Y})$) we denote the set of all polymorphisms (resp. all polymorphisms of arity $\ell$) of $\mathbb{Y}$. The core of a $\sigma$-structure $\mathbb{X}$—in symbols, $\operatorname{core}(\mathbb{X})$—is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$. It is easy to check that $\operatorname{core}(\mathbb{X})$ is unique up to isomorphism, and any endomorphism is an automorphism (see e.g. [HellNesetrilCore]).
The primitive positive (pp) fragment of first-order logic consists of the formulas that are built out of atomic formulas using conjunctions and existential quantifiers. If $\mathbb{A}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{A}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{A}$. If $R\subseteq A^{n}$ is a relation, we say that it is pp-definable over $\mathbb{A}$ if there exists a pp-formula defining it. A classical fact about pp-formulas is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{A}$, $(a_{1},\dots,a_{n})$ is in the relation defined by $\varphi$ in $\mathbb{A}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{A}$ that maps $x_{i}$ to $a_{i}$.
2.2. Quantum CSPs
Let $H$ be a non-trivial finite-dimensional Hilbert space. We say that a linear endomorphism $p$ of $H$ is an orthogonal projector (or simply a projector) if it is self-adjoint and idempotent (i.e., if $p^{2}=p=p^{*}$). We denote by $\operatorname{id}_{H}$ (or simply by $\operatorname{id}$ when the Hilbert space is clear) the identity projector over $H$. A projective measurement (PVM) over a Hilbert space $H$ and a set $B$ and is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$. A quantum function from $A$ to $B$ is a family of PVMs $(Q_{a,b})_{b\in B}$, one PVM for each $a\in A$. We shall denote this by the notation $Q\colon A\to_{q}B$ or $Q\colon A\to_{H}B$ if we want to emphasize the Hilbert space. We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B,b^{\prime}\in B^{\prime}$. Given a quantum function $Q:A\to_{q}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$. We say that $Q$ is contextual if it is not non-contextual. Note that each classical function $f:A\to B$ can be seen as a non-contextual quantum function.
Let $\mathbb{A},\mathbb{B}$ be relational structures with the same signature $\sigma$. A quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$, first defined at this level of generality in [abramsky2017quantum], is given by a non-trivial Hilbert space $H$ and a quantum function $Q\colon A\to_{q}B$ such that the following conditions hold:
  • ($\mathbf{QH_{1}}$)
    $\prod_{i\in[\operatorname{ar}(R)]}Q_{a_{i},b_{i}}=0\quad\forall R\in\sigma,\mathbf{a}\in R^{\mathbb{A}},\mathbf{b}\in B^{\operatorname{ar}(R)}\setminus R^{\mathbb{B}}$;
  • ($\mathbf{QH_{2}}$)
    $[Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0\quad\forall(a,a^{\prime})\in E(\operatorname{Gaif}(\mathbb{A})),(b,b^{\prime})\in B^{2}$.
Note: Condition $($$\mathbf{QH_{2}}$)) means that $Q$ corresponds to perfect oracular correlations. In Section 7 we shall extend our results to the non-oracular case. We also point out that the notion of a quantum homomorphism can be phrased equivalently as a representation of a certain finitely generated *-algebra [MastelSlofstra2024].
Definition 1 (Entangled CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{*}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{A}$.
2.3. Operations on quantum homomorphisms
We describe here certain constructions taking as input two quantum homomorphisms and giving another quantum homomorphism as the output.
Recall that if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\oplus Q^{\prime}$ is the operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$. We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}:X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}:X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$. Note that this is well defined as the set $\{Q_{x,a}\oplus Q^{\prime}_{x,a}\}_{a\in A}$ is a PVM over the Hilbert space $H\oplus H^{\prime}$. Indeed, the fact that its members are projectors onto $H\oplus H^{\prime}$ immediately follows from the definition. Moreover, it is easy to verify that the projector $\sum_{a\in A}(Q_{x,a}\oplus Q^{\prime}_{x,a})$ acts like the identity on both $H$ and $H^{\prime}$—and, hence, on the whole $H\oplus H^{\prime}$ by linearity.
Similarly, if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\otimes Q^{\prime}$ is the linear operator on $H\otimes H^{\prime}$ defined by letting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity. As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$. This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$.
Lemma 2.
Let $Q\colon\mathbb{X}\to_{H}\mathbb{A}$ and $Q^{\prime}\colon\mathbb{X}\to_{H^{\prime}}\mathbb{A}$ be quantum homomorphisms. Then $Q\oplus Q^{\prime}$ is a quantum homomorphism $\mathbb{X}\to_{H\oplus H^{\prime}}\mathbb{A}$.
Proof.
Let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$ and $(a_{1},\dots,a_{r})\not\in R^{\mathbb{A}}$. Then $R_{x_{1},a_{1}}\ldots R_{x_{r},a_{r}}=Q_{x_{1},a_{1}}\ldots Q_{x_{r},a_{r}}\oplus Q^{\prime}_{x_{1},a_{1}}\ldots Q^{\prime}_{x_{r},a_{r}}=0$, so $($$\mathbf{QH_{1}}$)) holds.
For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in\{1,\dots,r\}$, and $a,a^{\prime}\in A$. Then $R_{x_{i},a},R_{x_{j},a^{\prime}}]=[Q_{x_{i},a},Q_{x_{j},a^{\prime}}]\oplus[Q^{\prime}_{x_{i},a},Q^{\prime}_{x_{j},a^{\prime}}]=0$, so $($$\mathbf{QH_{2}}$)) holds. ∎
Observe that classical homomorphisms $h\colon A\to B$ can be identified with quantum homomorphisms $h\colon A\to_{H}B$ over a Hilbert space $H$ of dimension $1$. We will implicitly use this identification in the following. Now, if $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$). This is a simple consequence of the spectral theorem.
Definition 3.
Let $\mathbb{A},\mathbb{B}$ be two $\sigma$-structures. The quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ is the set $\oplus{\operatorname{Hom}(\mathbb{A},\mathbb{B})}$ of all quantum homomorphisms of the form $h_{1}\oplus\dots\oplus h_{d}$, where $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$.
[…]
(1) $Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}[h_{i}(a)=b]\langle x,e_{i}\rangle e_{i}$
for all $x\in H$, where $\mathbf{1}{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
Lemma 4.
Let $Q\colon\mathbb{X}\to_{H}\mathbb{A}$ and $Q^{\prime}\colon\mathbb{X}\to_{H^{\prime}}\mathbb{B}$ be quantum homomorphisms. Then $Q\otimes Q^{\prime}$ is a quantum homomorphism $\mathbb{X}\to_{H\otimes H^{\prime}}\mathbb{A}\times\mathbb{B}$.
Proof.
Let $R=Q\otimes Q^{\prime}$. Let $x\in X,a\in A,b\in B$. Then
$R_{x,(a,b)}(R_{x,(a,b)}(v\otimes v^{\prime}))$ $=R_{x,(a,b)}(Q_{x,a}(v)\otimes Q^{\prime}_{x,b}(v^{\prime}))$
$=Q_{x,a}(Q_{x,a}(v))\otimes Q^{\prime}_{x,b}(Q^{\prime}_{x,b}(v^{\prime}))$
$=Q_{x,a}(v)\otimes Q^{\prime}_{x,b}(v^{\prime})$
$=R_{x,(a,b)}(v\otimes v^{\prime})$
so that each $R_{x,(a,b)}$ is a projector. Moreover, if $x\in X$, then
$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}$
$=\sum_{a}Q_{x,a}\otimes(\sum_{b}Q_{x^{\prime},b})$
$=\sum_{a}Q_{x,a}\otimes\operatorname{id}_{H^{\prime}}$
$=(\sum_{a}Q_{x,a})\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}$
so $R$ is a quantum function. The verification that $($$\mathbf{QH_{1}}$)) holds is similar.
For $($$\mathbf{QH_{2}}$)), let $x,y$ appear together in a tuple in a relation of $\mathbb{X}$, and let $a,a^{\prime}\in A,b,b^{\prime}\in B$. Then
$R_{x,(a,b)}R_{y,(a^{\prime},b^{\prime})}$ $=(Q_{x,a}\otimes Q^{\prime}_{x,b})(Q_{y,a^{\prime}}\otimes Q^{\prime}_{y,b^{\prime}})$
$=(Q_{x,a}Q_{y,a^{\prime}})\otimes(Q^{\prime}_{x,b}Q^{\prime}_{y,b^{\prime}})$
$=(Q_{y,a^{\prime}}Q_{x,a})\otimes(Q^{\prime}_{y,b^{\prime}}Q^{\prime}_{x,b})$
$=R_{y,(a^{\prime},b^{\prime})}R_{x,(a,b)}$
so $($$\mathbf{QH_{2}}$)) holds. ∎
We shall also make use of the following composition operation for quantum homomorphisms.
Definition 5.
Let $Q\colon A\to_{H}B$ and $R\colon B\to_{H^{\prime}}C$ be quantum functions. We define
$R\bullet Q\colon A\to_{H\otimes H^{\prime}}C$
by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$.
Note that the sum in the definition is finite (even if $B$ is infinite), since $H$ has finite dimension and therefore only finitely many of the $Q_{a,b}$ are non-zero.
The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [abramsky2017quantum] using that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad. For the convenience of the reader, we give here a self-contained proof following [abramsky2017quantum].
Lemma 6 ([abramsky2017quantum]).
The composition of quantum homomorphisms $Q\colon\mathbb{A}\to_{H}\mathbb{B}$ and $R\colon\mathbb{B}\to_{H^{\prime}}\mathbb{C}$ is a quantum homomorphism $\mathbb{A}\to_{H\otimes H^{\prime}}\mathbb{C}$.
Proof.
We first verify that $S=R\bullet Q$ is a quantum function. For $a\in A,c\in C$, we get
$S_{a,c}S_{a,c}$ $=\sum_{b,b^{\prime}}(R_{b,c}\otimes Q_{a,b})(R_{b^{\prime},c}\otimes Q_{a,b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b,c}R_{b^{\prime},c})\otimes(Q_{a,b}Q_{a,b^{\prime}})$
$=\sum_{b}R_{b,c}\otimes Q_{a,b}$ (since $Q_{a,b}Q_{a,b^{\prime}}=0$ for $b\neq b^{\prime}$)
$=S_{a,c}.$
Similarly,
$S_{a,c}^{\top}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{\top}=\sum_{b}R_{b,c}^{\top}\otimes Q_{a,b}^{\top}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
so $S_{a,c}$ is a projector. We also have
$\sum_{c}S_{a,c}=\sum_{b,c}R_{b,c}\otimes Q_{a,b}=\sum_{b}\operatorname{id}\otimes Q_{a,b}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}.$
This means that $S$ is indeed a quantum function. Let us prove it is in fact a quantum homomorphism.
If $(a_{1},\dots,a_{r})\in R^{\mathbb{A}}$ and $(c_{1},\dots,c_{r})\not\in R^{\mathbb{C}}$ for some symbol $R\in\sigma$ of arity, say, $r$, then
$S_{a_{1},c_{1}}\dots S_{a_{r},c_{r}}$ $=\sum_{b_{1},\dots,b_{r}\in B}(R_{b_{1},c_{1}}\otimes Q_{a_{1},c_{1}})\dots(R_{b_{r},c_{r}}\otimes Q_{a_{r},b_{r}})$
$=\sum_{b_{1},\dots,b_{r}\in B}(R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}})\otimes(Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}).$
Note that for every $(b_{1},\dots,b_{r})$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)). Thus, the sum above evaluates to 0 and $($$\mathbf{QH_{1}}$)) holds.
If $a,a^{\prime}$ appear together in a tuple in a relation of $\mathbb{A}$ and $c,c^{\prime}\in C$, then
$S_{a,c}S_{a^{\prime},c^{\prime}}$ $=\sum_{b,b^{\prime}}(R_{b,c}\otimes Q_{a,b})(R_{b^{\prime},c^{\prime}}\otimes Q_{a^{\prime},b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b,c}R_{b^{\prime},c^{\prime}})\otimes(Q_{a,b}Q_{a^{\prime},b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b^{\prime},c^{\prime}}R_{b,c})\otimes(Q_{a^{\prime},b^{\prime}}Q_{a,b})$ by $($$\mathbf{QH_{2}}$))
$=S_{a^{\prime},c^{\prime}}S_{a,c}.$
so $($$\mathbf{QH_{2}}$)) holds. ∎
We will also make use of the following simple result.
Lemma 7.
Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{q}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
$\sum_{x\in X}Q_{x,y}=\operatorname{id}$
for each vertex $y$ of $\mathbb{X}$.
Proof.
Since $Q$ is non-contextual, we can write it as $Q=\bigoplus_{i\in[d]}h_{i}$ for some $d\in\mathbb{N}$ and some classical homomorphisms $h_{i}:\mathbb{X}\to\mathbb{X}$. By the definition of a core, each $h_{i}$ is an automorphism of $\mathbb{X}$. In particular, for each $i\in[d]$ there exists a unique $x_{i}$ such that $h_{i}(x_{i})=y$. It follows that
$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id}.$
3. Quantum polymorphisms and reductions
In this section, we show that reductions between entangled CSPs can be obtained via a quantised notion of pp-definitions that we call q-definitions. In order to prove that q-definitions exist, we make use of the notion of commutativity gadgets as defined in [Ji]. Then, we prove that the existence of commutativity gadgets is entirely determined by a closure property of the quantum polymorphisms of the given language. We begin by defining q-definitions.
Definition 8.
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
  • $(\mathbf{q_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
  • $(\mathbf{q_{2}})$
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$.
We can then define a notion of q-definability, as follows.
Definition 9.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-definable from $\mathbb{A}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
Observe, in ($(\mathbf{q_{1}})$), that a quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ is necessarily non-contextual, since all the vertices of $\mathbb{R}$ lie in a common constraint. (This will not be true anymore in the non-oracular setting, see Section 7.) Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages. The next result shows that the same is true for $\operatorname{CSP^{*}}$, provided that pp-definitions are replaced by the stronger q-definitions.
Proposition 10.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{B}$ is q-definable from $\mathbb{A}$. Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{*}}(\mathbb{A})$ in logspace.
We prove Proposition 10 later in this section. For now, we focus on the following problem: How to construct q-definitions guaranteeing that a given $\operatorname{CSP^{*}}$ can be reduced to a different one? It is not hard to show that every q-definition yields a classical pp-definition. Indeed, considering ($(\mathbf{q_{1}})$) for quantum homomorphisms with a Hilbert space of dimension 1 (which corresponds to classical homomorphisms), we find that the tuples in $R$ are in one-to-one correspondence with the restrictions of homomorphisms $\mathbb{G}\to\mathbb{A}$ to $\{g_{1},\dots,g_{r}\}$. However, not every pp-definition is a q-definition. It turns out that, in order for a pp-definition to ensure the existence of a q-definition, we need to pair it with an extra construction making sure that the process of replacing constraints with gadgets preserves the existence of quantum homomorphisms. This is the notion of commutativity gadgets we define below, generalising [Ji].
Definition 11.
Fix a $\sigma$-structure $\mathbb{A}$. A commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
  • $(\mathbf{c_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ extending $Q$.
  • $(\mathbf{c_{2}})$
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
Note that a commutativity gadget as in Definition 11 is not the same thing as a q-definition for $\mathbb{A}$ and $R=A^{2}$ in the sense of Definition 8.
Remark 12.
In [Zeman], an algebraic commutativity gadget is defined by replacing ($(\mathbf{c_{1}})$) by the following condition:
  • $(\mathbf{c_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
Note that this is a special case of $(\mathbf{c_{1}})$ and, thus, any commutativity gadget is an algebraic commutativity gadget. While this appears to be a strictly weaker requirement than the one in Definition 11, we shall prove below that the existence of an algebraic commutativity gadget in the sense of [Zeman] is equivalent to the existence of a commutativity gadget in our sense.
If $\mathbb{A}$ has a commutativity gadget, then we have a natural way of obtaining a sound reduction from $\operatorname{CSP^{*}}(\mathbb{B})$ to $\operatorname{CSP^{*}}(\mathbb{A})$ out of a classical gadget: For every pair of variables $x,y$ in a relation in an instance of $\operatorname{CSP^{*}}(\mathbb{B})$, glue to $x,y$ a copy of the commutativity gadget, identifying $x$ with $u$ and $y$ with $v$. The second condition in Definition 11 forces the associated projectors to commute in every quantum homomorphism to $\mathbb{A}$, and the first condition tells us that no additional constraint is forced upon $x,y$ by gluing the commutativity gadget. Formally, we prove the following result.
Proposition 13.
Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{B}$ is pp-definable from $\mathbb{A}$. Then $\mathbb{B}$ is q-definable from $\mathbb{A}$.
As a consequence of what we have discussed so far, the crucial step in order to quantise classical reductions from CSP theory is to build commutativity gadgets. The main result of this section classifies the languages for which commutativity gadgets exist, in terms of a property of their quantum polymorphisms, which we now define.
Definition 14.
A quantum polymorphism of arity $n\in\mathbb{N}$ for a structure $\mathbb{A}$ is a quantum homomorphism $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$.
We denote by $\operatorname{qPol}(\mathbb{A})$ the set of quantum polymorphisms of $\mathbb{A}$. Moreover, we define the quantum closure of $\operatorname{Pol}(\mathbb{A})$, denoted by $\operatorname{\oplus Pol}(\mathbb{A})$, as the union of the quantum closures of the set of $n$-ary polymorphisms $\operatorname{Hom}(\mathbb{A}^{n},\mathbb{A})$, for all $n\in\mathbb{N}$.
Remark 15.
Observe that $\operatorname{qPol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are both abstract minions (see the definition in [BBKO21, ciardo_quantum_minion]). Furthermore, it is not hard to check that $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are homomorphically equivalent, and $\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{A})$. In particular, it follows that, if $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, then we have a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{A})$.
The main result of this section is the following.
Theorem 16.
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
    $\mathbb{A}$ has a commutativity gadget.
  2. (2)
    $\mathbb{A}$ has an algebraic commutativity gadget.
  3. (3)
    $\mathbb{A}^{n}$ is an (algebraic) commutativity gadget for all large enough $n\in\mathbb{N}$.
  4. (4)
    $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
In the rest of this section, we provide the proofs of the results we stated.
Proof of Proposition 10.
For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$). Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{*}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$. For $R\in\rho$ of some arity $r$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$, and we glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{q}\mathbb{A}^{\prime}$. Then, ($(\mathbf{q_{1}})$) guarantees the existence of a quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$ such that
$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
for each $i\in[r]$ and each $a\in A$. Overall, this is a quantum homomorphism $\mathbb{Y}\to_{q}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\to_{q}\mathbb{A}$ is a quantum homomorphism. Consider the quantum function $Q^{\prime}\colon X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Observe that the restriction of $Q$ to $\mathbb{G}_{R}^{(\mathbf{x})}$ is a quantum homomorphism $\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$. By ($(\mathbf{q_{2}})$), this means that the assignment $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). In particular, recalling that $x_{i}$ and $g_{i}^{(\mathbf{x})}$ are identified in $\mathbb{Y}$, it follows that the PVMs $Q^{\prime}_{x_{i}}=Q_{g_{i}^{(\mathbf{x})}}$ and $Q^{\prime}_{x_{j}}=Q_{g_{j}^{(\mathbf{x})}}$ commute for each choice of $i,j\in[r]$. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(2) $\prod_{i\in[r]}Q^{\prime}_{x_{i},b_{i}}=0,$
which means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
Proof of Proposition 13.
Let $S=R^{\mathbb{B}}$ be a relation of $\mathbb{B}$ of some arity $r$, and let $\mathbb{G}$ be a classical gadget for $R^{\mathbb{B}}$ in $\mathbb{A}$, with distinguished vertices $g_{1},\dots,g_{r}$. Let $\mathbb{H}$ be a commutativity gadget for $\mathbb{A}$, with distinguished vertices $u,v$. We let $\mathbb{G}^{\prime}$ be the structure that initially consists of the disjoint union of $\mathbb{G}$ and a copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$ for each pair of distinct vertices $x,y$ in $G$, and in which we identify each such pair $x,y$ with the vertices $u,v$ from the copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$.
We show that $\mathbb{G}^{\prime}$ (with distinguished vertices $g_{1},\dots,g_{r}$) is a q-definition for $S$ in $\mathbb{A}$. Let $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ be a quantum homomorphism. In particular, $Q$ is non-contextual and can therefore be written as $h_{1}\oplus\dots\oplus h_{d}$ for some homomorphisms $h_{k}\colon\mathbb{R}\to\mathbb{A}_{S}$. Since $\mathbb{G}$ is a gadget for $S$ in $\mathbb{A}$, it follows that there exist homomorphisms $f_{k}\colon\mathbb{G}\to\mathbb{A}$ such that $f_{k}(g_{i})=h_{k}(i)$ for all $i\in[r]$. Then $Q^{\prime}=\bigoplus_{k\in[d]}f_{k}$ is a non-contextual quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$. In particular, $Q^{\prime}_{x}$ and $Q^{\prime}_{y}$ commute for all $x,y\in G$, and it follows from ($(\mathbf{c_{1}})$) that for all $x,y$ there exists a quantum homomorphism $Q^{(x,y)}\colon\mathbb{H}\to_{q}\mathbb{A}$ such that $Q^{(x,y)}_{u}=Q^{\prime}_{x}$ and $Q^{(x,y)}_{v}=Q^{\prime}_{y}$. We thus obtained a quantum homomorphism $\mathbb{G}^{\prime}\to_{q}\mathbb{A}$, which proves ($(\mathbf{q_{1}})$).
Let now $Q\colon\mathbb{G}^{\prime}\to_{q}\mathbb{A}$ be a quantum homomorphism. For each $x,y\in G$, we have by ($(\mathbf{c_{2}})$) that $Q_{x}$ and $Q_{y}$ commute. Thus, $Q$ restricts to a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ that is non-contextual and can thus be written as $h_{1}\oplus\dots\oplus h_{d}$, each $h_{k}$ being a classical homomorphism $\mathbb{G}\to\mathbb{A}$. Since $\mathbb{G}$ is a classical gadget for $R$, we obtain that $(h_{k}(g_{1}),\dots,h_{k}(g_{r}))\in R^{\mathbb{B}}$ for all $k\in[d]$. Thus, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}=\bigoplus_{k}(h_{k})_{g_{i},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$. ∎
Next, we show that q-definitions are compatible with quantum homomorphisms.
Proposition 17.
Let $\mathbb{A},\mathbb{B}$ be structures, $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively. Then every quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{q}(B;R_{B})$.
Proof.
Let $g_{1},\dots,g_{r}$ be the distinguished vertices of $\mathbb{G}$.
We prove $($$\mathbf{QH_{1}}$)). Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$. By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in\{1,\dots,r\}$. Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{B}$. By ($(\mathbf{q_{2}})$), we have that the quantum function defined by $Q^{\prime\prime}_{i,b_{i}}:=S_{g_{i},b_{i}}$ for all $i\in\{1,\dots,r\}$ is a quantum homomorphism $Q^{\prime\prime}\colon\mathbb{R}\to(B,R_{B})$. Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$. Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$. Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$.
We now turn to $($$\mathbf{QH_{2}}$)). Let $a_{1},a_{2}$ appear together in a tuple in $R_{A}$ at positions $i$ and $j$ and let $b_{1},b_{2}\in B$ be arbitrary. Let $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ be such that $Q^{\prime}_{g_{i},a_{1}}=Q^{\prime}_{g_{j},a_{2}}=\operatorname{id}$, whose existence is asserted by ($(\mathbf{q_{1}})$). Again, let $S\colon\mathbb{G}\to_{q}\mathbb{B}$ be the composition $Q\bullet Q^{\prime}$ and $Q^{\prime\prime}\colon\mathbb{R}\to_{q}(B,R_{B})$ be the quantum homomorphism given by ($(\mathbf{q_{2}})$). We get $0=[R_{i,b_{1}},R_{j,b_{2}}]=[S_{g_{i},b_{1}},S_{g_{j},b_{2}}]=[Q_{a_{1},b_{1}}\otimes\operatorname{id},Q_{a_{2},b_{2}}\otimes\operatorname{id}]=[Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]\otimes\operatorname{id}$ and therefore $Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]=0$. ∎
In the rest of this section, we give a proof of Theorem 16. We start by observing the following fact.
Proposition 18.
A quantum polymorphism of a structure $\mathbb{A}$ is non-contextual if and only if it belongs to $\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
This is a simple consequence of the spectral theorem. ∎
Definition 19.
Let $\mathbb{A}$ be a structure. The pairs $(a_{1},b_{1}),\dots,(a_{n},b_{n})\in A^{2}$ are called generators of $A^{2}$ in $\operatorname{Pol}(\mathbb{A})$ if for every pair $(c,d)\in A^{2}$, there exists $g\in\operatorname{Pol}(\mathbb{A})$ of arity $n$ such that $g(a_{1},\dots,a_{n})=c$ and $g(b_{1},\dots,b_{n})=d$.
Note that, if $A$ is finite (as we always assume), $A^{2}$ admits a family of $n=|A|^{2}$ generators. Indeed, we can simply let $g$ be the $i$-th projection $\mathbb{A}^{n}\to\mathbb{A}$, with $i\in[n]$ being the index of the pair $(c,d)$. However, $n$ could be smaller for certain $\mathbb{A}$. We point out that there are many cases of infinite structures $\mathbb{A}$ where $A^{2}$ is finitely generated under $\operatorname{Pol}(\mathbb{A})$; this is for example the case for the class of $\omega$-categorical structures, which has been widely studied in the context of classical constraint satisfaction problems.
Proposition 20.
Let $\mathbb{A}$ be a structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$. Suppose also that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Then $\mathbb{A}^{n}$ is a commutativity gadget for $\mathbb{A}$.
Proof.
Let $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ be generators for $A^{2}$ under $\operatorname{Pol}(\mathbb{A})$. Let $\mathbb{G}$ be $\mathbb{A}^{n}$ and the two distinguished elements be $\mathbf{a}=(a_{1},\dots,a_{n})$ and $\mathbf{b}=(b_{1},\dots,b_{n})$.
Every quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function. Using that $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ are generators for $A^{2}$, we know that, for each $i$, there exists $g_{i}\in\operatorname{Pol}(\mathbb{A})$ such that $g_{i}(\mathbf{a})=h_{i}(\mathbf{a})$ and $g_{i}(\mathbf{b})=h_{i}(\mathbf{b})$. Then $Q$ extends to a quantum homomorphism $Q^{\prime}:=g_{1}\oplus\cdots\oplus g_{d}$, thus proving ($(\mathbf{c_{1}})$). ∎
We now prove that the existence of a commutativity gadget guarantees that all quantum polymorphisms are classical. In fact, we prove that the same fact holds even for the algebraic commutativity gadgets from Remark 12.
Proposition 21.
Let $\mathbb{A}$ be a structure. If $\mathbb{A}$ has an algebraic commutativity gadget, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Let $\mathbb{G}$ be an algebraic commutativity gadget with distinguished elements $u,v\in G$. Let $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$. We want to show that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, which is enough to conclude by Proposition 18.
Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that
$W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}.$
Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{q}\mathbb{A}^{n}$ given by the tensor product
$W=\bigotimes_{i\in[n]}W^{(i)}$
and observe that $W_{u,\mathbf{a}}=W_{v,\mathbf{b}}=\operatorname{id}$. Take now the composition
$U=Q\bullet W:\mathbb{G}\to_{q}\mathbb{A}.$
Using ($(\mathbf{c_{2}})$), we observe that $U$ satisfies $U_{u,c},U_{v,d}]=0$. However, expanding the composition as per Definition 5, we have
$U_{u,c}=\sum_{\mathbf{t}\in A^{n}}Q_{\mathbf{t},c}\otimes W_{u,\mathbf{t}}=Q_{\mathbf{a},c}\otimes\operatorname{id}$
and, similarly, $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Thus,
$Q_{\mathbf{a},c},Q_{\mathbf{b},d}]\otimes\operatorname{id}=[Q_{\mathbf{a},c}\otimes\operatorname{id},Q_{\mathbf{b},d}\otimes\operatorname{id}]=[U_{u,c},U_{v,d}]=0,$
which implies that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we are done. ∎
Proof of Theorem 16.
The result is a direct consequence of Propositions 20 and 21 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 12). ∎
4. Warm-up: Quantum polymorphisms of cliques
In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques. In particular, we prove the following.
Theorem 22.
For all $m\geq 3$, $\operatorname{qPol}(\mathbb{K}_{m})=\operatorname{\oplus Pol}(\mathbb{K}_{m})$.
Proof.
The proof goes by induction on $m\geq 3$.
For the case of $m=3$, it follows from [Ji, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 16 implies the result.
Take now $m\geq 4$. Let $Q\colon(\mathbb{K}_{m})^{n}\to_{q}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$. Pick $\mathbf{u},\mathbf{v}\in[m]^{n}$ and $i,j\in[m]$. Our goal is to show that $Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=0$.
Choose $\mathbf{w}\in[m]^{n}$ in such a way that $(\mathbf{u},\mathbf{w})$ and $(\mathbf{v},\mathbf{w})$ are both edges of $(\mathbb{K}_{m})^{n}$. (Note that this is possible as long as $m\geq 3$.) Choose also $k\in[m]\setminus\{i,j\}$. We first claim that
(3) $[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$
This is clearly the case if $Q_{\mathbf{w},k}=0$, so assume this is not the case. Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in\{1,\dots,n\}$, namely the subgraph induced by the neighborhood of $\mathbf{w}$. Note that $\mathbb{X}$ is isomorphic to $(\mathbb{K}_{m-1})^{n}$.
Define $R\colon\mathbb{X}\to_{q}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$. We check that this is a quantum homomorphism. Every $R_{\mathbf{z},\ell}$ is a projector since $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell}]=0$. For $\mathbf{z}\in X$, we have $\sum_{\ell}R_{\mathbf{z},\ell}=\sum_{\ell}Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}=Q_{\mathbf{w},k}(\sum_{\ell}Q_{\mathbf{z},\ell})=Q_{\mathbf{w},k}$, which is indeed the identity function on $Q_{\mathbf{w},k}(H)$ since $Q_{\mathbf{w},k}$ is a projector. Thus, $R$ is a quantum function. If $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$, then $\mathbf{z},\mathbf{z}^{\prime},\mathbf{w}$ form a triangle in $(\mathbb{K}_{m})^{n}$ and therefore the projectors $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell},Q_{\mathbf{z}^{\prime},\ell^{\prime}}$ pairwise commute for $\ell,\ell^{\prime}$ distinct from $k$. Thus, $($$\mathbf{QH_{2}}$)) holds. Finally, if $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$ and $\ell\neq k$, then $R_{\mathbf{z},\ell}R_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{w},k}Q_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{z}^{\prime},\ell}=0$. Thus, $($$\mathbf{QH_{1}}$)) holds.
By the inductive hypothesis, we have $\operatorname{qPol}(\mathbb{K}_{m-1})=\operatorname{\oplus Pol}(\mathbb{K}_{m-1})$. Therefore, $R_{\mathbf{z},\ell},R_{\mathbf{z}^{\prime},\ell^{\prime}}]=0$ for all $\mathbf{z},\mathbf{z}^{\prime}\in X$ and all $\ell,\ell^{\prime}\in[m]\setminus\{k\}$. For $\mathbf{z}=\mathbf{u}$ and $\mathbf{z}^{\prime}=\mathbf{v}$, this gives the claimed eq. 3.
We can now already conclude by noting that
$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}],$
that all the terms with $k=\ell$ in the sum vanish by eq. 3, and that all the terms with $k\neq\ell$ in the sum vanish, since
$[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}]$ $=Q_{\mathbf{u},i}Q_{\mathbf{w},k}Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}-Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}Q_{\mathbf{u},i}Q_{\mathbf{w},k}$
$=Q_{\mathbf{u},i}Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}Q_{\mathbf{v},j}-Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}Q_{\mathbf{u},i}$
$=0,$
where the second equality is due to the fact that $Q_{\mathbf{w},k}$ and $Q_{\mathbf{w},\ell}$ commute with $Q_{\mathbf{u},i}$ and $Q_{\mathbf{v},j}$, and the third holds since $Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}=0=Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}$. ∎
Note that $\mathbb{K}_{2}$ has quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$. Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{q}\mathbb{K}_{2}$ by the following matrices:
  • $Q_{(0,0),0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},Q_{(0,0),1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$
  • $Q_{(1,1),0}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},Q_{(1,1),0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$
  • $Q_{(1,0),0}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix},Q_{(1,0),1}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$
  • $Q_{(0,1),0}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},Q_{(0,1),1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$.
It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$. By Theorem 16, we obtain that $\mathbb{K}_{2}$ does not have a commutativity gadget.
Proposition 23.
Let $m\geq 3$. Then $\mathbb{K}_{m}$ has a commutativity gadget. In fact, for every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
Proof.
We apply Proposition 20. The assumptions on $\mathbf{u},\mathbf{v}$ imply that $(u_{1},v_{1}),\dots,(u_{n},v_{n})$ generate $m]^{2}$. Indeed, any pair $(c,c)$ is the image of $(u_{i},v_{i})$ under a permutation (which is in particular a unary polymorphism of $\mathbb{K}_{m}$), while any pair $(c,d)$ with $c\neq d$ is the image of $(u_{j},v_{j})$ under a permutation. ∎
A consequence of Proposition 23 is the undecidability of the entangled $m$-coloring problem for $m\geq 3$. The case of $m=3$ was proved in [culf_re-completeness_2025], and the case for $m>3$ was recently and independently proved in [Zeman, Proposition 5.2, Lemma 4.6]. We include a simple proof here using commutativity gadgets for the convenience of the reader.
Corollary 24.
$\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
Proof.
We shall reduce from $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, whose undecidability was proved in [Ji, culf2024re]. Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, which we can assume does not contain isolated vertices. Let $\mathbb{Y}$ be the graph obtained by adjoining m3m-3 new vertices $y_{4},\dots,y_{m}$ to $\mathbb{X}$ and such that, whenever $x$ is a vertex in $\mathbb{X}$, then $\{x,y_{4},\dots,y_{m}\}$ forms a clique in $\mathbb{Y}$. Finally, adjoin a commutativity gadget between every $y_{i}$ and every other vertex in $\mathbb{Y}$. Such a commutativity gadget exists by Proposition 23. We let $\mathbb{Z}$ be the resulting structure.
If $\mathbb{X}$ has a quantum homomorphism $Q$ to $\mathbb{K}_{3}$, then $Q^{\prime}\colon\mathbb{Y}\to_{q}\mathbb{K}_{m}$ defined by $Q^{\prime}_{y_{i},i}=\operatorname{id}$ for all $i\in\{4,\dots,m\}$ and $Q_{x,j}=Q^{\prime}_{x,j}$ for all $x\in X$ is a quantum homomorphism. This $Q^{\prime}$ can be extended to a quantum homomorphism $\mathbb{Z}\to\mathbb{K}_{m}$ by ($(\mathbf{c_{1}})$).
Conversely, suppose that there exists a quantum homomorphism $Q\colon\mathbb{Z}\to_{q}\mathbb{K}_{m}$. By ($(\mathbf{c_{2}})$), it follows that for every $x\in X$, the restriction of $Q$ to $\{x,y_{2},\dots,y_{4}\}$ is non-contextual. Thus, this restriction can be written as $h_{1}\oplus\dots\oplus h_{d}$ for some classical homomorphisms $h_{1},\dots,h_{d}$. Up to composing the $h_{i}$ with an automorphism of $\mathbb{K}_{m}$, we can assume without loss of generality that $h_{i}(y_{j})=j$ for all $i\in\{1,\dots,d\}$ and $j\in\{4,\dots,m\}$. Thus, for each $i\in\{1,\dots,d\}$ and $x\in X$, we have that $h_{i}(x)\in\{1,2,3\}$. It follows that $\sum_{i=1}^{3}Q_{x,i}=\operatorname{id}$ and that $Q$ restricts to a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{K}_{3}$. ∎
5. Intermezzo: Contextuality and bifurcations
In this section, we present a general result on the emergence of a certain non-orthogonality structure in any contextual quantum homomorphism. While in the current work we shall only use this result to analyse quantum polymorphisms of odd cycles (see Section 6), we believe this result will be useful in a broader setting.
$Q_{x_{0}}$$Q_{x_{1}}$$Q_{x_{2}}$$a_{0}$$a_{0}^{\prime}$$a_{1}$$a_{2}^{\prime}$$a_{2}$
Figure 1. A bifurcation of length $2$. Solid lines mean non-orthogonality.
The high-level idea is that, whenever some non-commutativity occurs in a quantum homomorphism $Q:\mathbb{X}\to_{q}\mathbb{Y}$, a bifurcation must appear in the product $X\times A$, in a sense that we formalise below. For simplicity, we restrict the analysis to binary structures (i.e., digraphs), although the same argument applies to quantum homomorphisms between arbitrary structures.
Definition 25.
Let $\mathbb{X},\mathbb{A}$ be digraphs, let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer. A bifurcation for $Q$ is a set
$V=\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}\subseteq X\times A$
such that
  • $x_{0},x_{1},\dots,x_{\omega}$ is a path in $\mathbb{X}$;
  • $a_{0}\neq a_{0}^{\prime}$ and $a_{\omega}\neq a_{\omega}^{\prime}$;
  • $Q_{\alpha,\beta}Q_{\delta,\gamma}\neq 0$ for each $((\alpha,\beta),(\delta,\gamma))\in V^{2}\setminus\{((x_{0},a_{0}^{\prime}),(x_{\omega},a_{\omega}^{\prime})),((x_{\omega},a_{\omega}^{\prime}),(x_{0},a_{0}^{\prime}))\}$ such that $\alpha\neq\delta$. We call $\omega$ the length of the bifurcation.
Recall that $\mathbb{X}^{u}$ denotes the undirected graph underlying a digraph $\mathbb{X}$.
Theorem 26.
Let $\mathbb{X}$ and $\mathbb{A}$ be digraphs such that $\mathbb{X}^{u}$ is connected, and let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism. Then there exists a bifurcation for $Q$ whose length is at most the diameter of $\mathbb{X}^{u}$.
We will need the following result from [manvcinska2016quantum], of which we give an alternative proof making use of the machinary of Section 3.
Proposition 27 ([manvcinska2016quantum]).
Let $Q\colon\mathbb{X}\to_{q}\mathbb{Y}$ be a quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$. If $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$, then $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$.
Proof.
Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two extremities of this path. Since this structure is a tree, Proposition 47 implies that $(\mathbb{G},u,v)$ is a q-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively. It follows from Proposition 17 that $Q$ is a quantum homomorphism $(X,R_{X})\to_{q}(Y,R_{Y})$. By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)). ∎
We shall repeatedly use a simple fact about projectors, stated below without proof.
Fact 28.
Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if and only if $PQP=PQ=QP$.
Proof of Theorem 26.
Let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism over some Hilbert space $H$ of dimension $h$. Let $S$ be the set of pairs $(x,x^{\prime})\in X^{2}$ for which there exist $(a,a^{\prime})\in A^{2}$ such that $Q_{x,a},Q_{x^{\prime},a^{\prime}}]\neq 0$. Since $Q$ is contextual, $S$ is non-empty. Choose $(x,x^{\prime})\in S$ in a way that $\operatorname{dist}(x,x^{\prime})$ is minimum over all pairs in $S$ (where the distance is computed in $\mathbb{X}^{u}$), and call such distance $\omega$. Note that $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)). Let $x=x_{0},x_{1},\dots,x_{\omega}=x^{\prime}$ be a shortest oriented path in $\mathbb{X}$ connecting $x$ and $x^{\prime}$. By the minimality of $\operatorname{dist}(x,x^{\prime})$, observe that the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\beta}}$ commute for each $\alpha\leq\beta\in\{0,\dots,\omega\}$ such that $\alpha\neq 0$ or $\beta\neq\omega$. In particular, we can choose an o.n. basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$. On the other hand, since $Q_{x_{0}}$ and $Q_{x_{\omega}}$ do not commute, there exist $a_{0},a_{\omega}\in A$ such that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
Take the set $Z=\{i\in[h]:Q_{x_{\omega},b}e_{i}\in\{0,e_{i}\}\}$, and consider the projectors $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$ and $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$. Observe that $P_{1}P_{2}=0$ and $P_{1}+P_{2}=\operatorname{id}$. Moreover, we have
(4) $[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$
Indeed, $P_{1}$ is diagonal in the same basis $\mathcal{E}$ that diagonalises each $Q_{x_{\alpha}}$, so it commutes with all projectors in such PVMs. Furthermore,
$Q_{x_{\omega},a_{\omega}}P_{1}=\sum_{i\in Z}Q_{x_{\omega},a_{\omega}}e_{i}e_{i}^{*}=\sum_{i\in Z}e_{i}e_{i}^{*}Q_{x_{\omega},a_{\omega}}=P_{1}Q_{x_{\omega},a_{\omega}}$
since, for $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$. Hence, $P_{1}$ also commutes with $Q_{x_{\omega},a_{\omega}}$. Finally, $P_{1},M]+[P_{2},M]=[\operatorname{id},M]=0$ for each $M$ because of the bilinearity of the commutator, so eq. 4 holds. In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors. Observe now that
$[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]$ $=[Q_{x_{0},a_{0}}P_{1}+Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}+Q_{x_{\omega},a_{\omega}}P_{2}]$
$=[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{1}]+[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]$
(5) $+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$
Combining eq. 4 and 28, we find that the second summand in the right-hand side of section 5 is
$[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]=[P_{1}Q_{x_{0},a_{0}}P_{1},P_{2}Q_{x_{\omega},a_{\omega}}P_{2}]=0$
since $P_{1}P_{2}=0$. Similarly, we obtain $Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]=0$. Consider now the first summand in the right-hand side of section 5. Both the projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$. Hence, they commute, too. We deduce that
(6) $[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$
We now claim that the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ pairwise commute. The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is trivial, since $\mathcal{E}$ diagonalises both. Moreover,
$Q_{x_{\alpha},d}P_{2}Q_{x_{\omega},a_{\omega}}P_{2}=Q_{x_{\alpha},d}Q_{x_{\omega},a_{\omega}}P_{2}=Q_{x_{\omega},a_{\omega}}Q_{x_{\alpha},d}P_{2}=Q_{x_{\omega},a_{\omega}}P_{2}Q_{x_{\alpha},d}P_{2},$
where the first and third equalities come from eq. 4 and 28, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by the minimality of $\operatorname{dist}(x,x^{\prime})$. So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some o.n. basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
For the next step, it will be useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$. For $m\in M$, consider also the set $A_{m}=\{i\in[h]:{\langle e_{i},f_{m}\rangle}\neq 0\}$. Observe that, for any $m\in M$,
$Q_{x_{0},a_{0}}P_{2}f_{m}f_{m}^{*}$ $=Q_{x_{0},a_{0}}P_{2}\left(\sum_{i\in[h]}e_{i}e_{i}^{*}\right)f_{m}f_{m}^{*}\left(\sum_{j\in[h]}e_{j}e_{j}^{*}\right)=\sum_{i,j\in[h]}Q_{x_{0},a_{0}}P_{2}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}$
$=\sum_{\begin{subarray}{c}i\in L\cap A_{m}\\ j\in A_{m}\end{subarray}}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}$
and
$f_{m}f_{m}^{*}Q_{x_{0},a_{0}}P_{2}$ $=\left(\sum_{i\in[h]}e_{i}e_{i}^{*}\right)f_{m}f_{m}^{*}\left(\sum_{j\in[h]}e_{j}e_{j}^{*}\right)Q_{x_{0},a_{0}}P_{2}=\sum_{i,j\in[h]}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}Q_{x_{0},a_{0}}P_{2}$
$=\sum_{\begin{subarray}{c}i\in A_{m}\\ j\in L\cap A_{m}\end{subarray}}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}.$
Hence, if $A_{m}\subseteq L$ or if $L\cap A_{m}=\emptyset$, it holds that $Q_{x_{0},a_{0}}P_{2},f_{m}f_{m}^{*}]=0$. On the other hand, we know from eq. 6 that
$0\neq[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=\sum_{m\in M}[Q_{x_{0},a_{0}}P_{2},f_{m}f_{m}^{*}],$
whence we deduce that there must exist some $\bar{m}\in M$ for which $A_{\bar{m}}\not\subseteq L$ and $A_{\bar{m}}\not\subseteq[h]\setminus L$.
We now choose $\xi_{1}\in A_{\bar{m}}\cap L$ and $\xi_{2}\in A_{\bar{m}}\cap([h]\setminus L)$. We claim that
(7) $\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$
Indeed, suppose for the sake of contradiction that $\xi_{i}\in Z$ for some $i\in[2]$. This would imply that $P_{2}e_{\xi_{i}}=0$, whence in turn we could deduce that
$\sum_{m\in M}f_{m}f_{m}^{*}e_{\xi_{i}}=Q_{x_{\omega},a_{\omega}}P_{2}e_{\xi_{i}}=0.$
It would follow that ${\langle e_{\xi_{i}},f_{m}\rangle}=0$ for each $m\in M$. In particular, $\xi_{i}\not\in A_{\bar{m}}$, which is a contradiction. Hence, (7) is true.
Let now $a_{0}^{\prime}\in A$ be such that $Q_{x_{0},a_{0}^{\prime}}e_{\xi_{2}}=e_{\xi_{2}}$. Notice that there exists exactly one such $a_{0}^{\prime}$, since $Q_{x_{0}}$ is a PVM diagonalised by $\mathcal{E}$. We claim that $a_{0}^{\prime}\neq a_{0}$. Suppose, for the sake of contradiction, that $a_{0}^{\prime}=a_{0}$. Then, it would hold that $Q_{x_{0},a_{0}}e_{\xi_{2}}=e_{\xi_{2}}$. On the other hand, we have
(8) $0=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}e_{\xi_{2}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{2}}=P_{2}Q_{x_{0},a_{0}}e_{\xi_{2}}=P_{2}e_{\xi_{2}}=e_{\xi_{2}},$
where the first equality is true since $\xi_{2}\not\in L$, the second is true by definition of $L$, the third by eq. 4, the fourth by the observation above, and the fifth by eq. 7. This is a contradiction, so the claimed disequality $a_{0}^{\prime}\neq a_{0}$ must hold.
Fix now some $\alpha\in[\omega-1]$, and recall that the PVM $Q_{x_{\alpha}}$ is diagonalised by $\mathcal{E}$. Hence, there exists exactly one element $c_{1}\in A$ such that $Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=e_{\xi_{1}}$, and exactly one element $c_{2}\in A$ such that $Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=e_{\xi_{2}}$. We claim that $c_{1}=c_{2}$. Suppose this is not the case. It is time to remember that we chose $\mathcal{F}$ in such a way that it diagonalises both $Q_{x_{\omega},a_{\omega}}P_{2}$ and all the matrices $Q_{x_{\alpha},d}P_{2}$ for $d\in A$. Hence, we can write $Q_{x_{\alpha},c_{1}}P_{2}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}$ and $Q_{x_{\alpha},c_{2}}P_{2}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}$ for some $V_{1},V_{2}\subseteq[h]$. Since $Q_{x_{\alpha}}$ is a PVM, we have $Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}=0$. Using eq. 4 and 28, we deduce that
$0=Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}P_{2}=Q_{x_{\alpha},c_{1}}P_{2}Q_{x_{\alpha},c_{2}}P_{2},$
so $V_{1}\cap V_{2}=\emptyset$. On the other hand, using eq. 7, we have
$e_{\xi_{1}}$ $=Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=Q_{x_{\alpha},c_{1}}P_{2}e_{\xi_{1}}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}e_{\xi_{1}}\quad\mbox{and}$
$e_{\xi_{2}}$ $=Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=Q_{x_{\alpha},c_{2}}P_{2}e_{\xi_{2}}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}e_{\xi_{2}}.$
Since we are assuming that $\{\xi_{1},\xi_{2}\}\subseteq A_{\bar{m}}$, we must have $f_{\bar{m}}^{*}e_{\xi_{1}}\neq 0$ and $f_{\bar{m}}^{*}e_{\xi_{2}}\neq 0$. But this means that $\bar{m}\in V_{1}\cap V_{2}$, a contradiction. This proves the claimed equality $c_{1}=c_{2}$. Call such element $a_{\alpha}$.
Observe now that
$Q_{x_{0},a_{0}}e_{\xi_{1}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{1}}=e_{\xi_{1}},$
where the first equality is due to eq. 7 and the second to the fact that $\xi_{1}\in L$ and, for each $\ell\in L$, $e_{\ell}$ is an eigenvector of $Q_{x_{0},a_{0}}P_{2}$ with eigenvalue $1$. It follows that, for each $i\in[\omega-1]$,
$Q_{x_{0},a_{0}}Q_{x_{i},a_{i}}e_{\xi_{1}}=Q_{x_{0},a_{0}}e_{\xi_{1}}=e_{\xi_{1}},$
which means that $Q_{x_{0},a_{0}}Q_{x_{i},a_{i}}\neq 0$. Similarly,
$Q_{x_{0},a_{0}^{\prime}}Q_{x_{i},a_{i}}e_{\xi_{2}}=Q_{x_{0},a_{0}^{\prime}}e_{\xi_{2}}=e_{\xi_{2}},$
so $Q_{x_{0},a_{0}^{\prime}}Q_{x_{i},a_{i}}\neq 0$. Moreover, clearly, $Q_{x_{i},a_{i}}Q_{x_{j},a_{j}}\neq 0$ for each $i\neq j\in[\omega-1]$ since $e_{\xi_{1}}$ is an eigenvector of both with eigenvalue $1$.
Observe now that
$Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=f_{\bar{m}}$
since $\bar{m}\in M$ and, for each $m\in M$, $f_{m}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}P_{2}$ with eigenvalue $1$. Hence, for each $i\in[\omega-1]$,
$e_{\xi_{1}}^{*}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=e_{\xi_{1}}^{*}f_{\bar{m}}\neq 0$
since $\xi_{1}\in A_{\bar{m}}$. In particular, this means that $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}\neq 0$. Analogously,
$e_{\xi_{2}}^{*}Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=e_{\xi_{2}}^{*}f_{\bar{m}}\neq 0$
since $\xi_{2}\in A_{\bar{m}}$. Hence, $Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}\neq 0$. Note also that the $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}=0$ is trivial from the fact that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
To conclude, we claim that there exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
(9) $\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$
Note that this would be enough to finish the proof. Indeed, it follows from the minimality of $\operatorname{dist}(x,x^{\prime})$ that (9) guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $0\leq i\leq\omega-1$.
Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}$ $=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\operatorname{id}=\sum_{d\in A}\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},d}$
$=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}.$
But then, we would have
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}e_{\xi_{1}}=e_{\xi_{1}}$
so, in particular,
$\|\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|=1.$
However, we know from eq. 7 that $\xi_{1}\not\in Z$, which by the definition of $Z$ implies that $Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\neq e_{\xi_{1}}$. Since $Q_{x_{\omega},a_{\omega}}$ is a projector, this means that $\|Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|<1$. Since
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}$
is also a projector, it follows that
$\|\prod_{i\in[\omega-1]}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|<1,$
a contradiction. Hence, the claim is true.
Putting all together, we conclude that the vertices
$\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}$
form a bifurcation for $Q$ of length $\omega$, and the proof is complete. ∎
[…]
Figure 2. A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
6. Quantum polymorphisms of odd cycles
The goal of this section is to prove the following.
Theorem 29.
For each odd $m\geq 3$, it holds that $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$.
We begin with a simple consequence of Theorem 26, which handles the case of quantum polymorphisms of arity 11—i.e., quantum endomorphisms.
Corollary 30.
$\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})=\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ for each odd $m\geq 3$.
Proof.
Suppose, for the sake of contradiction, that $\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})\neq\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$, and let $Q$ be a contextual quantum homomorphism $\mathbb{C}_{m}\to_{q}\mathbb{C}_{m}$. Applying Theorem 26, we deduce that there exists a bifurcation for $Q$ of length $\omega\leq\frac{m-1}{2}$. Consider, in particular, the initial elements $(x_{0},a_{0}),(x_{0},a_{0}^{\prime}),(x_{1},a_{1}),(x_{2},a_{2})$ of such bifurcation. By Definition 25 we know that $Q_{x_{0},a_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{0},a^{\prime}_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$, $Q_{x_{0},a_{0}}Q_{x_{2},a_{2}}\neq 0$, and $Q_{x_{0},a_{0}^{\prime}}Q_{x_{2},a_{2}}\neq 0$. Using $($$\mathbf{QH_{1}}$)), we deduce that $(a_{0},a_{1}),(a_{0}^{\prime},a_{1})$, and $(a_{1},a_{2})$ are all edges of $\mathbb{C}_{m}$. Note also that $x_{0}$ and $x_{2}$ lie on an (m2)(m-2)-walk in $\mathbb{X}$. Since no walk of such length connects a vertex to itself in $\mathbb{C}_{m}$Proposition 27 ensures that $a_{0}\neq a_{2}$ and $a_{0}^{\prime}\neq a_{2}$. Finally $a_{0}\neq a_{0}^{\prime}$ again by Definition 25. We deduce that $a_{1}$ has degree at least three in $\mathbb{C}_{m}$, a contradiction. ∎
We point out that a different proof of Corollary 30 was obtained in [Zeman] via a result from [banica2007quantum] on the quantum symmetry of vertex-transitive graphs. In the following, we fix an odd integer $m\in\mathbb{N}$. We start with the following observation.
Lemma 31.
Let $Q$ be a quantum endomorphism of $\mathbb{C}_{m}$ and let $x,y,z$ be a $2$-walk in $\mathbb{C}_{m}$. Then $Q_{x,a}Q_{y,b}Q_{z,c}=Q_{y,b}Q_{z,c}$ for each $a,b,c\in\mathbb{Z}_{m}$.
Proof.
If $x=z$ the result is trivial so we may assume that $x\neq z$. We know from Corollary 30 that there exists $d\in\mathbb{N}$ and $h_{1},\dots h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$. Using eq. 1, we can write
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)$
$=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
Let $i\in[d]$ give a non-trivial contribution to the sum above. Since $x,y,z$ form a path in $\mathbb{C}_{m}$, any choice of $(b,c)\in\mathbb{Z}_{m}^{2}$ for which $h_{i}(y)=b$ and $h_{i}(z)=c$ gives precisely one choice of $a\in\mathbb{Z}_{m}$ such that $h_{i}(x)=a$. It follows that
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}=Q_{y,b}Q_{z,c},$
as needed. ∎
Next, we focus on quantum polymorphisms of arity at least two. Henceforth in this section, we fix some arity $n\geq 2$. Furthermore, whenever a plus or minus sign is applied to an element of $\mathbb{Z}_{m}$ or to a tuple of such elements, it shall be meant modulo $m$.
Lemma 32.
Fix two vertices $\mathbf{x},\mathbf{y}$ of $\mathbb{C}_{m}^{n}$ such that $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For any $a\in\mathbb{Z}_{m}$, it holds that
$[Q_{\mathbf{x},a},Q_{\mathbf{y},a}]=0.$
Proof.
In this proof, we shall abbreviate $Q_{\mathbf{a},a}$ to $Q_{\mathbf{a}}$ for each $\mathbf{a}\in[m]^{n}$. Let $\mathbf{x}\to\mathbf{z}\to\mathbf{y}$ be a $2$-path connecting $\mathbf{x}$ and $\mathbf{y}$ in $\mathbb{C}_{m}^{n}$. Let also $\mathbf{u}=\mathbf{z}-\mathbf{x}$ and $\mathbf{v}=\mathbf{y}-\mathbf{z}$, and notice that $\mathbf{u},\mathbf{v}\in\{\pm 1\}^{n}$. Observe that the vertices $\mathbf{y},\mathbf{y}+\mathbf{v},\mathbf{y}+2\mathbf{v},\dots,\mathbf{y}+(m-1)\mathbf{v}$ induce a $\mathbb{C}_{m}$ within $\mathbb{C}_{m}^{n}$. The restriction of $Q$ to such vertices is a quantum endomorphism of $\mathbb{C}_{m}$. Therefore, combining Corollary 30 and Lemma 7, we obtain that
$\sum_{i\in\mathbb{Z}_{m}}Q_{\mathbf{y}+i\mathbf{v}}=\operatorname{id}.$
Consider the two paths
$\mathbf{x}\to\mathbf{x}+\mathbf{u}\to\mathbf{y}\to\mathbf{y}+\mathbf{v}\to\mathbf{y}+2\mathbf{v}\to\dots\to\mathbf{y}+(m-4)\mathbf{v},$
$\mathbf{x}\to\mathbf{x}+\mathbf{u}\to\mathbf{y}-2\mathbf{v}\to\mathbf{y}-3\mathbf{v}\to\dots\to\mathbf{y}-(m-2)\mathbf{v},$
and notice that both have length m2m-2. Since there is no closed walk in $\mathbb{C}_{m}$ of odd length less than $m$, it follows from Proposition 27 that $Q_{\mathbf{x}}Q_{\mathbf{y}+i\mathbf{v}}=0$ for each $i\in\mathbb{Z}_{m}\setminus\{0,-2\}$. Therefore,
$0$ $=Q_{\mathbf{y}+\mathbf{v}}Q_{\mathbf{x}}=\left(\operatorname{id}-\sum_{i\in\mathbb{Z}_{m}\setminus\{1\}}Q_{\mathbf{y}+i\mathbf{v}}\right)Q_{\mathbf{x}}=Q_{\mathbf{x}}-(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}.$
Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are Hermitian, it follows that
$[Q_{\mathbf{x}},Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}}]=0.$
Consider now the automorphism $\xi$ of $\mathbb{C}_{m}^{n}$ that acts as the reflection around $x_{\alpha}$ for each coordinate $\alpha\in[n]$ for which $u_{\alpha}=v_{\alpha}$, and acts as the identity for all other coordinates. Note that both $\mathbf{x}$ and $\mathbf{x}+\mathbf{u}-\mathbf{v}=\mathbf{y}-2\mathbf{v}$ are fixed points of $\xi$, while $\xi(\mathbf{y})=2\mathbf{x}-\mathbf{y}$. Since the previous part of the argument is invariant under (classical) automorphisms of $\mathbb{C}_{m}^{n}$, we deduce that
$[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}}]=0.$
By the bilinearity of the commutator, it follows that
$[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}].$
We now apply the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
$[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}]$ $=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{2\mathbf{x}-\mathbf{y}}]=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{4\mathbf{x}-3\mathbf{y}}]=\dots$
$=[Q_{m\mathbf{x}-(m-1)\mathbf{y}},Q_{(m+1)\mathbf{x}-m\mathbf{y}}]=[Q_{\mathbf{y}},Q_{\mathbf{x}}],$
whence we conclude that $Q_{\mathbf{x}},Q_{\mathbf{y}}]=0$, as required. ∎
Lemma 33.
Fix two vertices $\mathbf{x},\mathbf{y}$ of $\mathbb{C}_{m}^{n}$ such that $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For any $a,b\in\mathbb{Z}_{m}$, it holds that
$[Q_{\mathbf{x},a},Q_{\mathbf{y},b}]=0.$
Proof.
We can assume without loss of generality that $b=a\pm 2$ $\mod m$, as otherwise $Q_{\mathbf{x},a}Q_{\mathbf{y},b}=0$ by Proposition 27. Up to symmetry, we can further suppose that $b=a+2$ $\mod m$. Let $\mathbf{u}$ and $\mathbf{v}$ be as in the proof of Lemma 32, and observe that
$Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}$ $=Q_{\mathbf{x},a}\left(\sum_{b\in\mathbb{Z}_{m}}Q_{\mathbf{y}-\mathbf{v},b}\right)\left(\sum_{c\in\mathbb{Z}_{m}}Q_{\mathbf{y}-2\mathbf{v},c}\right)Q_{\mathbf{y},a+2}$
$=\sum_{b,c\in\mathbb{Z}_{m}}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},b}Q_{\mathbf{y}-2\mathbf{v},c}Q_{\mathbf{y},a+2}.$
Since $Q_{\mathbf{y}-\mathbf{v},b},Q_{\mathbf{y}-2\mathbf{v},c}]=0$, for a product in the summation above to be non-zero, it must be the case that $b$ is a neighbour of both $a$ and $a+2$; i.e., $b=a+1$. Consider now $Q_{\mathbf{y}-2\mathbf{v},c}$. For the product to be non-zero, since $\mathbf{y}-2\mathbf{v}$ is a $\mathbb{C}_{m}^{n}$-neighbour of $\mathbf{y}-\mathbf{v}$, it must hold that $c$ is a $\mathbb{C}_{m}$-neighbour of $b$; i.e., $c=a$ or $c=a+2$. Suppose $Q_{\mathbf{y}-2\mathbf{v},a+2}\neq 0$. We know from Lemma 32 that $Q_{\mathbf{y}-2\mathbf{v},a+2},Q_{\mathbf{y},a+2}]=0$ and, hence, the two projectors can be diagonalised simultaneously. Hence, if it was the case that $Q_{\mathbf{y}-2\mathbf{v},a+2}Q_{\mathbf{y},a+2}\neq 0$, the ranges of $Q_{\mathbf{y}-2\mathbf{v},a+2}$ and of $Q_{\mathbf{y},a+2}$ would have non-trivial intersection. But this would imply that $\sum_{i\in\mathbb{Z}_{m}}Q_{\mathbf{y}+i\mathbf{v},a+2}\neq\operatorname{id}$, contradicting Corollary 30 and Lemma 7. It follows that the only choice for $c$ giving a non-trivial contribution is $c=a$, so
$Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}=Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}.$
By a similar argument, we have
$Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}$ $=Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}.$
Hence, we obtain
$Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}$ $=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}$
$=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}$
$=Q_{\mathbf{y},a+2}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{x},a}$
(where we are using that $Q_{\mathbf{x},a},Q_{\mathbf{y}-\mathbf{v},a+1}]=0$ as $\mathbf{x}\sim\mathbf{y}-\mathbf{v}$ and $Q_{\mathbf{x},a},Q_{\mathbf{y}-2\mathbf{v},a}]=0$ by Lemma 32)
$=Q_{\mathbf{y},a+2}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{x},a}$
$=(Q_{\mathbf{x},a}Q_{\mathbf{y},a+2})^{*}=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}.$
In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is a Hermitian matrix, so
$[Q_{\mathbf{x},a},Q_{\mathbf{y},a+2}]=0,$
as needed. ∎
We can now finalise the proof of the main result of this section.
Proof of Theorem 29.
Let $\mathbf{x}$ and $\mathbf{y}$ be two vertices of $\mathbb{C}_{m}^{n}$ and let $a,b$ be two vertices of $\mathbb{C}_{m}$. We need to show that $Q_{\mathbf{x},a},Q_{\mathbf{y},b}]=0$. We shall use induction on $\delta=\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})$. If $\delta\in\{0,1\}$, the result is clear from the definition of quantum homomorphisms. If $\delta=2$, the result is Lemma 33. Let then $\delta\geq 3$, and suppose that the result holds for all pairs of vertices at mutual distance at most $\delta-1$. Let $\mathbf{x},\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{\delta-1},\mathbf{y}$ be a shortest path from $\mathbf{x}$ to $\mathbf{y}$ in $\mathbb{C}_{m}^{n}$. Since $\delta\geq 3$, we can assume without loss of generality that the vectors $\mathbf{z}_{1}-\mathbf{x}$ and $\mathbf{z}_{2}-\mathbf{z}_{1}$ are equal. [^2] Note that this is false in general when $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For instance, consider the case that $n=2$, $m\geq 5$, $\mathbf{x}=(0,0)$, and $\mathbf{y}=(0,2)$. We now expand the product $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ as follows:
$Q_{\mathbf{x},a},Q_{\mathbf{y},b}$ $=Q_{\mathbf{x},a}\left(\sum_{c_{1}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{1},c_{1}}\right)\left(\sum_{c_{2}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{2},c_{2}}\right)\dots\left(\sum_{c_{\delta-1}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{\delta-1},c_{\delta-1}}\right)Q_{\mathbf{y},b}$
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}Q_{\mathbf{y},b}.$
Note now that $\operatorname{dist}(\mathbf{z}_{i},\mathbf{y})\leq\delta-1$ for each $i\in[\delta-1]$. Hence, the inductive hypothesis allows us to freely permute the corresponding projectors in the expression above, thus yielding
(10) $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
Multiplying both terms by $Q_{\mathbf{x},a}$ on the right and further rearranging gives
$Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}Q_{\mathbf{x},a}$
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}$
Observe now that $\mathbf{x}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ lie on a common $m$-cycle within $\mathbb{C}_{m}^{n}$ since $\mathbf{z}_{1}-\mathbf{x}=\mathbf{z}_{2}-\mathbf{z}_{1}$. By Lemma 31, it holds that $Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}=Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}$ for each $c_{1},c_{2}\in\mathbb{Z}_{m}$. We deduce that
$Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}=Q_{\mathbf{x},a}Q_{\mathbf{y},b},$
where the second equality is eq. 10. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required.
Corollary 34.
Let $m\geq 3$ be an odd natural number. Then $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$. In particular, $\mathbb{C}_{m}$ has a commutativity gadget and $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable.
Proof.
$\mathbb{C}_{m}$ has a commutativity gadget by Theorem 29 and Theorem 16.
Note that $\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a simple path of length m2m-2. Since $\mathbb{C}_{m}$ has a commutativity gadget, it follows that $\mathbb{K}_{m}$ is q-definable in $\mathbb{C}_{m}$ by Proposition 13. By Proposition 10, $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ reduces to $\operatorname{CSP^{*}}(\mathbb{C}_{m})$, and $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable by Corollary 24. Thus, $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable. ∎
We conclude with the proof that quantum homomorphisms between different odd cycles can be contextual.
Proposition 35.
There is a quantum homomorphism $\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
Proof.
Consider for this proof $\mathbb{C}_{5}$ to have domain $\{0,1,2,5,6\}$ with edges between adjacent numbers, and $\mathbb{C}_{7}$ to have domain $[7]$. Consider the two classical endomorphisms $h,g\colon\mathbb{C}_{5}\to\mathbb{C}_{5}$ such that $h(2)=6,h(5)=0$ and $g(5)=0,g(2)=1$. Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space. Extend $Q$ by setting $Q_{3,a}=Q_{5,a}$ for $a\in\{0,1,2,5,6\}$. Choose a basis in which all the projectors are diagonal, and in this basis define $Q_{4,1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ and $Q_{4,6}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$. Then $Q\colon\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ and $Q_{4,0},Q_{2,1}]\neq 0$. ∎
7. The non-oracular setting
In this section, we consider the non-oracular version of the results in Section 3. A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{qno}}\mathbb{B}$, is a quantum function $Q\colon A\to_{q}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)).
The following is the non-oracular version of the entangled CSP.
Definition 36 (Non-oracular entangled CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{**}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{qno}}\mathbb{A}$.
We now define quantum polymorphisms in the non-oracular setting.
Definition 37.
Let $\mathbb{A}$ be a relational structure. A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We denote by $\operatorname{qPol^{no}}(\mathbb{A})$ the set of non-oracular quantum polymorphisms of $\mathbb{A}$.
We will obtain a quantisation of the Galois connection between classical polymorphisms and relations. The following is the main result of this section.
Theorem 38.
Let $\mathbb{A},\mathbb{B}$ be structures on potentially different signatures such that
$\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B}).$
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
We begin by defining the non-oracular version of q-definitions.
Definition 39.
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A non-oracular q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
  1. $(\mathbf{noq_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
  2. $(\mathbf{noq_{2}})$
    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}_{S}$.
Note that, in the first item of the definition above, $Q$ is required to be oracular (i.e., to satisfy $($$\mathbf{QH_{2}}$))), while $Q^{\prime}$ need not.
Definition 40.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-no-definable from $\mathbb{A}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
In the following, it shall be useful to observe that Lemma 2, Lemma 7Lemma 4, and Lemma 6 also hold in the non-oracular setting. Furthermore, the definition of quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ (Definition 3) does not change in the non-oracular setting.
The following result is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
Theorem 41.
Let $\mathbb{A}$ be a finite structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$. Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{A}_{S})$.
Proof.
Henceforth in this proof, we shall denote $\mathbb{A}_{S}$ by $\mathbb{D}$ for the sake of readability.
  • $(\Rightarrow)$
    Let $\mathbb{G}$ be a non-oracular q-definition for $\mathbb{A}$ and $S$. This means that every quantum homomorphism $\mathbb{R}\to_{q}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{D}$.
    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We want to prove that $Q\in\operatorname{qPol^{no}}(\mathbb{D})$. Note that $Q$ is a quantum function $A^{n}\to_{q}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
    Fix then such a tuple $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})\in R^{\mathbb{D}^{n}}$, and choose a tuple $\mathbf{c}\in A^{r}\setminus S$. For every $i\in[n]$, consider the deterministic homomorphism $f_{i}:\mathbb{R}\to\mathbb{D}$ defined by $f_{i}(j)=b^{(j)}_{i}$ for each $j\in[r]$. In particular, $f_{i}$ is a quantum homomorphism. Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $f_{i}$. Consider now the tensor product
    $W=\bigotimes_{i\in[n]}W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}^{n},$
    and compose it with $Q$ to obtain $Q\bullet W:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$. Fix $\ell\in[r]$ and $a\in A$. Expanding the composition and the tensor product yields
    $(Q\bullet W)_{g_{\ell},a}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes W_{g_{\ell},\mathbf{a}}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes\bigotimes_{i\in[n]}W^{(i)}_{g_{\ell},a_{i}}$.
    We know from ($(\mathbf{noq_{1}})$) that $W^{(i)}_{g_{\ell},a_{i}}=(f_{i})_{\ell,a_{i}}$. Since $f_{i}$ is deterministic, this means that in order for the $\mathbf{a}$-th summand in the sum above to be non-zero it must be that $a_{i}=f_{i}(\ell)=b_{i}^{(\ell)}$ for each $i\in[n]$; i.e., $\mathbf{a}=\mathbf{b}^{(\ell)}$. Hence,
    $(Q\bullet W)_{g_{\ell},a}=Q_{\mathbf{b}^{(\ell)},a}\otimes\operatorname{id}.$
    On the other hand, ($(\mathbf{noq_{2}})$) implies that the assignment $(\ell,a)\mapsto(Q\bullet W)_{g_{\ell},a}$ yields a non-oracular quantum homomorphism from $\mathbb{R}$ to $\mathbb{D}$. Since we are assuming that $\mathbf{c}\not\in S=R^{\mathbb{D}}$, we deduce that
    $0=\prod_{\ell\in[r]}(Q\bullet W)_{g_{\ell},c_{\ell}}=\prod_{\ell\in[r]}(Q_{\mathbf{b}^{(\ell)},a}\otimes\operatorname{id})=\left(\prod_{\ell\in[r]}Q_{\mathbf{b}^{(\ell)},a}\right)\otimes\operatorname{id}.$
    Therefore, we find $\prod_{\ell\in[r]}Q_{\mathbf{b}^{(\ell)},a}=0$. It follows that $Q$ is a non-oracular quantum homomorphism from $\mathbb{D}^{n}$ to $\mathbb{D}$; i.e., an element of $\operatorname{qPol^{no}}(\mathbb{D})$, as desired.
  • $(\Leftarrow)$
    Suppose now that $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{D})$. Let us enumerate the tuples in $S$ as $\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}$, where $n=|S|$. We will show that $\mathbb{A}^{n}$ is a non-oracular quantum gadget for $\mathbb{A}$ and $S$, with distinguished vertices $g_{1},\dots,g_{r}$ defined by $g_{\ell}=(a^{(1)}_{\ell},\dots,a^{(n)}_{\ell})$ for $\ell\in[r]$. We show that the two conditions in Items $(\mathbf{noq_{1}})$ and $(\mathbf{noq_{2}})$ hold.
    Let $Q\colon\mathbb{R}\to_{q}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$. Since $Q$ satisfies $($$\mathbf{QH_{2}}$)), $Q$ must be classical, so by the spectral theorem we can decompose it as $Q=\bigoplus_{i\in[d]}h_{i}$, where $h_{i}\colon\mathbb{R}\to\mathbb{D}$ is a deterministic homomorphism for each $i\in[d]$. By the definition of $\mathbb{R}$, it must be that $(h_{i}(1),\dots,h_{i}(r))\in S$ for each $i\in[d]$. In other words, there exists a function $f:[d]\to[n]$ satisfying
    (11) $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$
    for each $i\in[d]$. Consider now the deterministic dictator polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$. Define
    $P=\bigoplus_{i\in[d]}\pi^{n}_{f(i)}.$
    Note that $P$ is a classical (a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
    $P_{g_{\ell},a}=\bigoplus_{i\in[d]}(\pi^{n}_{f(i)})_{g_{\ell},a}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ \pi^{n}_{f(i)}(g_{\ell})=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ a^{f(i)}_{\ell}=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ h_{i}(\ell)=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\end{subarray}}(h_{i})_{\ell,a}=Q_{\ell,a}$
    (where the fourth equality follows from Equation 11). Hence, $P$ witnesses that ($(\mathbf{noq_{1}})$) holds.
    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$. Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuples $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures. Since $Q$ is a non-oracular quantum polymorphism of $\mathbb{A}$, by assumption it is also an $n$-ary non-oracular quantum polymorphism of $\mathbb{D}$. Hence, for each $\mathbf{c}\in A^{r}\setminus S$, we must have
    $0=\prod_{\ell\in[r]}Q_{g_{\ell},c_{\ell}}=\prod_{\ell\in[r]}Q^{\prime}_{\ell,c_{\ell}},$
    which means that $Q^{\prime}$ is a non-oracular quantum homomorphism from $\mathbb{R}$ to $\mathbb{D}$. This shows that ($(\mathbf{noq_{2}})$) is also satisfied, and concludes the proof of the theorem.∎
The following result is an immediate consequence of Theorem 41.
Corollary 42.
Two structures $\mathbb{A},\mathbb{B}$ satisfy $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ if and only if $\mathbb{B}$ is q-no-definable from $\mathbb{A}$.
Unlike in the oracular setting, non-oracular q-gadgets are not strong enough to guarantee reductions. To that end, they need to be strengthened via a non-oracular commutativity gadget, which we now define.
Definition 43.
Fix a $\sigma$-structure $\mathbb{A}$. A non-oracular commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
  1. $(\mathbf{noc_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $Q$.
  2. $(\mathbf{noc_{2}})$
    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
We now give the non-oracular analogue of Theorem 16. Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [Zeman], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 43 with the weaker condition stated below:
  1. $(\mathbf{noc_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
Theorem 44.
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
    $\mathbb{A}$ has a non-oracular commutativity gadget.
  2. (2)
    $\mathbb{A}$ has an algebraic non-oracular commutativity gadget.
  3. (3)
    $\mathbb{A}^{n}$ is an (algebraic) non-oracular commutativity gadget for all large enough $n\in\mathbb{N}$.
  4. (4)
    $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
We prove the following implications.
  • $(4.\Rightarrow 3.)$
    The proof follows from exactly the same argument as for Proposition 20.
  • $(3.\Rightarrow 1.)$
    Obvious.
  • $(1.\Rightarrow 2.)$
    This is clear, since each non-oracular commutativity gadget is in particular an algebraic non-oracular commutativity gadget (via the same argument as in Remark 12).
  • $(2.\Rightarrow 4.)$
    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$. We proceed as in Proposition 21. We first build a non-oracular quantum polymorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$). Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ as the composition of $Q$ and $W$. On the one hand, ($(\mathbf{noq_{2}})$) guarantees that $U_{u,c},U_{v,d}]=0$. On the other hand, expanding the composition yields $U_{u,c}=Q_{\mathbf{a},c}\otimes\operatorname{id}$ and $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Combining the two facts, we conclude that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we deduce that $Q$ is classical.∎
The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP^{*}}$s to $\operatorname{CSP^{**}}$s.
Proposition 45.
Let $\mathbb{A}$ be a $\sigma$-structure and let $\mathbb{B}$ be a $\rho$-structure having the same domain (but potentially distinct signatures $\sigma$ and $\rho$), and suppose that
  • $\mathbb{A}$ has a non-oracular commutativity gadget, and
  • for every $R\in\rho$, there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$.
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
Proof.
For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$). Let also $\mathbb{H}$ be a non-oracular commutativity gadget for $\mathbb{A}$ with its distinguished vertices $u,v$.
Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. Recall that $\operatorname{Gaif}(\mathbb{X})$ denotes the Gaifman digraph of $\mathbb{X}$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{**}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, as well as a copy $\mathbb{H}^{(x,y)}$ of $\mathbb{H}$ for each edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. For $R\in\rho$ of some arity $r\in\mathbb{N}$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$. Similarly, we denote by $u^{(x,y)}$ (resp. $v^{(x,y)}$) the copy of $u$ (resp. $v$) in $\mathbb{H}^{(x,y)}$. For every edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$, glue $u^{(x,y)}$ with $x$ and glue $v^{(x,y)}$ with $y$. For each $R\in\rho$ of some arity $r$, each $\mathbf{x}\in R^{\mathbb{X}}$, and each $i\in[r]$, glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to\mathbb{A}^{\prime}$. (Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{qno}}\mathbb{A}$ such that
$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
for each $i\in[r]$ and each $a\in A$. Fix now an edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. Using again the fact that $Q$ is oracular, we know that $Q_{x,a},Q_{y,b}]=0$ for every $a,b\in A$. Hence, we can apply ($(\mathbf{noc_{1}})$) to find a non-oracular quantum homomorphism
$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{qno}}\mathbb{A}$
extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$ is a non-oracular quantum homomorphism. Consider the quantum function $Q^{\prime}:X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is an (oracular) quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Take $i,j\in[r]$. We claim that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute. This is easily true if $x_{i}=x_{j}$ (as in this case the PVMs coincide), so suppose that $x_{i}\neq x_{j}$. In this case, $(x_{i},x_{j})$ is an edge in $\operatorname{Gaif}(\mathbb{X})$. Note now that, by construction of $\mathbb{Y}$, the vertices $x_{i}$, $u^{(x_{i},x_{j})}$, and $g_{i}^{(\mathbf{x})}$ are identified in $Y$. Similarly, the vertices $x_{j}$, $v^{(x_{i},x_{j})}$, and $g_{j}^{(\mathbf{x})}$ are identified in $Y$. Since $Q$ induces a non-oracular quantum homomorphism over $\mathbb{H}^{(x_{i},x_{j})}$, we deduce from ($(\mathbf{noc_{2}})$) that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute also in this case. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(12) $\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$
(Note that for this to be true we do not need $U$ to be oracular.) This means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum (oracular) homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
We can finally complete the proof of the main result of this section.
Proof of Theorem 38.
Since $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})$Theorem 44 guarantees that $\mathbb{A}$ has a non-oracular commutativity gadget. Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Corollary 42 that all the relations of $\mathbb{B}$ have a q-no-definition in $\mathbb{A}$. Note also that $\mathbb{A}$ and $\mathbb{B}$ have the same domain since $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$. Hence, the result directly follows from Proposition 45. ∎
We now show that tree gadgets can be used to produce q-no-definitions. A relational structure $\mathbb{X}$ is a tree if for every sequence of tuples $\mathbf{t}^{1},\dots,\mathbf{t}^{m}$ that all belong to (potentially different) relations $R_{i}$ of $\mathbb{X}$ with arity $r_{i}$, the total number of elements appearing in $\mathbf{t}^{1},\dots,\mathbf{t}^{m}$ is at least $1+\sum_{i=1}^{m}(r_{i}-1)$.
Lemma 46.
Let $\mathbb{T}$ be a tree structure, and let $x,y$ be elements of $\mathbb{T}$. Let $Q\colon\mathbb{T}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. If $Q_{x,a}Q_{y,b}\neq 0$, there exists a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ such that $h(x)=a$ and $h(y)=b$.
Proof.
We prove this by induction on the size of $\mathbb{T}$. Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in\{1,\dots,r\}$. For the argument below, we can assume without loss of generality that $i<j$ up to swapping $x$ and $y$. Then we have
$Q_{x,a}Q_{y,b}$ $=\sum_{c_{1},\dots,c_{r}\in A:c_{i}=a,c_{j}=b}\prod_{k=1}^{r}Q_{z_{k},c_{k}}$
$=\sum_{(c_{1},\dots,c_{r})\in R^{\mathbb{A}}:c_{i}=a,c_{j}=b}\prod_{k=1}^{r}Q_{z_{k},c_{k}}$ by $($$\mathbf{QH_{1}}$))
and since $Q_{x,a}Q_{y,b}\neq 0$, it must be that there exists $(c_{1},\dots,c_{r})\in R^{\mathbb{A}}$ such that $c_{i}=a$ and $c_{j}=b$. We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in\{1,\dots,r\}$.
Otherwise, $\mathbb{T}$ consists of the union of some trees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$ that share a single vertex $z$. Since $\sum_{c}Q_{x,a}Q_{z,c}Q_{z,c}Q_{y,b}=Q_{x,a}(\sum_{c}Q_{z,c})Q_{y,b}=Q_{x,a}Q_{y,b}\neq 0$, there exists a $c$ such that $Q_{x,a}Q_{z,c}\neq 0$ and $Q_{z,c}Q_{y,b}\neq 0$. By induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp. $y\in\mathbb{T}_{i}$) then $h_{i}(x)=a$ (resp. $h_{i}(y)=b$). The union of these homomorphisms gives a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ with $h(x)=a$ and $h(y)=b$. ∎
Proposition 47.
Let $\mathbb{A}$ be a structure, and let $R\subseteq A^{2}$ be a relation with a classical gadget $\mathbb{G}$ that is a tree. Then $\mathbb{G}$ is a non-oracular q-definition for $R$.
Proof.
Suppose that $Q\colon\mathbb{R}\to_{q}(A;R)$ is a quantum homomorphism. Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $\{1,\dots,r\}\to A$ such that $(h_{i}(1),h_{i}(2))\in R$. Thus, each $h_{i}$ can be extended to a homomorphism $g_{i}\colon\mathbb{G}\to\mathbb{A}$ since $\mathbb{G}$ is a classical gadget for $R$. It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ satisfying the conclusion of ($(\mathbf{noq_{1}})$).
To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. Let $x_{1},x_{2}$ be the distinguished vertices of $\mathbb{G}$. Suppose that $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$. Since $\mathbb{G}$ is a tree, there exists by Lemma 46 a homomorphism $h\colon\mathbb{G}\to\mathbb{A}$ such that $h(x_{i})=a_{i}$ for $i\in\{1,2\}$. Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}(A;R)$. ∎
We conclude this section by showing that also the non-oracular version of the CSP parameterised by odd cycles is undecidable. The result is proved in same way as its oracular counterpart.
Corollary 48.
For every odd $m\geq 3$, $\operatorname{CSP^{**}}(\mathbb{C}_{m})$ is undecidable.
Proof.
Simply observe that the proof of Theorem 29 does not make use of the fact that quantum polymorphisms of $\mathbb{C}_{m}$ are oracular. Hence, the same argument establishes that $\operatorname{qPol^{no}}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$ for each odd $m\geq 3$. Then, the result follows from Theorem 38 applied to $\mathbb{A}=\mathbb{B}=\mathbb{C}_{m}$ and Corollary 34. ∎
8. Quantum polymorphisms of Boolean structures
In this section, we provide some results regarding the quantum polymorphisms of Boolean structures. If a quantum homomorphism $Q$ has range $\{0,1\}$, then $Q_{x,0}$ is determined by $Q_{x,1}$ since $Q_{x,0}=\operatorname{id}-Q_{x,1}$. Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $\{1,\dots,n\}$. Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ iff $i\in S$. In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{q}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
8.1. The case of structures without a majority polymorphism
In this subsection we show that any Boolean relational structure which is not invariant under the majority polymorphism has a commutativity gadget. This specific statement has previously been proven in [culf_re-completeness_2025, Section 7], our contribution is to simplify the proof by using the characterization in Theorem 44 to assert the non-existence of commutativity gadgets through the study of quantum polymorphisms. This allows us to make use of standard CSP methods in the proof.
We focus our study on the following class of Boolean relational structures:
Definition 49.
Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x\oplus_{2}t\mid x\in R_{1/k}\})$, where
  1. (1)
    $R_{1/k}=\{x\in\{0,1\}^{k}\mid\sum_{i=1}^{k}x_{i}=1\}$ is the relation of 11-in-$k$ SAT and,
  2. (2)
    $\oplus_{2}$ is componentwise addition modulo 2.
The relation of $\mathbb{O}_{t}$ is named $R_{t}$.
One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any 3 distinct tuples in the relation yields $t\not\in R_{t}$, witnessing the non-preservation.
We show that if we translate 11-in-$k$ by a tuple $t\in R_{1/k}$, then $\operatorname{\oplus Pol}(\mathbb{O}_{t})=\operatorname{qPol^{no}}(\mathbb{O}_{t})$. First consider the case where $k=3$ and $t=(1,0,0)$:
Proposition 50.
For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$, the following properties hold:
  1. (1)
    If $S\cap T=\emptyset$, then $Q_{S}Q_{T}=0$.
  2. (2)
    If $S\cap T=\emptyset$, then $Q_{S\cup T}=Q_{S\cup T}(Q_{S}+Q_{T})$.
  3. (3)
    $Q_{S}=Q_{S\cup T}Q_{S}$ for all $S$
  4. (4)
    $Q_{S}=\sum_{i\in S}Q_{\{i\}}$.
Proof.
Consider $S,T\subseteq[n]$ with $S\cap T=\emptyset$ and let $R:=S\cup T$. Then $(r,s,t)$ is in $R_{(1,0,0)}$, which gives the following identities by property $($$\mathbf{QH_{1}}$)) of $Q$:
  • $Q_{S\cup T}Q_{S}Q_{T}=0$, since $(1,1,1)$ is not in $R_{(1,0,0)}$,
  • $(\operatorname{id}-Q_{S\cup T})Q_{S}Q_{T}=0$, since $(0,1,1)$ is not in $R_{(1,0,0)}$,
  • $Q_{S\cup T}(\operatorname{id}-Q_{S})(\operatorname{id}-Q_{T})=0$, since $(1,0,0)$ is not in $R_{(1,0,0)}$.
  • $(\operatorname{id}-Q_{S\cup T})Q_{S}(\operatorname{id}-Q_{T})=0$, since $(0,1,0)$ is not in $R_{(1,0,0)}$.
We can add the first two identities to get $Q_{S}Q_{t}=0$ so, (1) holds. Moreover, adding the first and third identities yields (2).
To derive $Q_{S}=Q_{S\cup T}Q_{S}$ we add the first and fourth identities to get $Q_{S}=Q_{S\cup T}Q_{S}+Q_{S}Q_{T}$. Application of (1) then yields (3).
Finally, from the equalities
$Q_{S\cup T}=Q_{S\cup T}(Q_{S}+Q_{T})=Q_{S\cup T}Q_{S}+Q_{S\cup T}Q_{T}=Q_{S}+Q_{T}$
we can derive (4) by inductively building any $Q_{S}$ from sums of the individual elements of $S$. ∎
In particular, it follows that in every non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$ and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
Lemma 51.
For any natural number $k\geq 3$, the translation $\mathbb{O}_{(1,0,\dots,0)}$ of 11-in-$k$ by $(1,0,\dots,0)$ has the property that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})=\operatorname{\oplus Pol}(\mathbb{O}_{(1,0,\dots,0)})$.
Proof.
Since $(0,0,\dots,0)=(1,0,\dots,0)\oplus_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
$a=0\Leftrightarrow R_{(1,0,\dots,0)}(a,\dots,a)$
This relation is preserved by every non-oracular quantum polymorphism of $R_{(1,0,\dots,0)}$ since,
$Q_{\mathbf{a},1}=Q_{\mathbf{a},1}\dots Q_{\mathbf{a},1}=0$
for every $\mathbf{a}\in\{0,1\}^{n}$ and $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to\mathbb{O}_{(1,0,\dots,0)}$. We then proceed to define $R_{(1,0,0)}$ in $(\{0,1\},R_{(1,0,\dots,0)},0)$ by
$R_{(1,0,0)}(x,y,z)\Leftrightarrow R_{(1,0,\dots,0)}(x,y,z,0,\dots,0)$
We prove that $R_{(1,0,0)}$ is preserved by every non-oracular $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to\mathbb{O}_{(1,0,\dots,0)}$. Assume we have tuples $\mathbf{a},\mathbf{b},\mathbf{c}\in\{0,1\}^{n}$ with $(\mathbf{a},\mathbf{b},\mathbf{c})$ in the $n$th direct power of $R_{(1,0,0)}$ and variables $x,y,z\in\{0,1\}$ such that $(x,y,z)\notin R_{(1,0,0)}$. Then $(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{0},\dots,\mathbf{0})$ is in the $n$th power $R_{(1,0,\dots,0)}$ and $(x,y,z,0,\dots,0)$ is not an element of $R_{(1,0,\dots,0)}$
Now assume we have some non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to_{q}\mathbb{O}_{(1,0,\dots,0)}$ and consider:
$Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}$ $=Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}\sum_{w\in\{0,1\}}Q_{\mathbf{0},w}$
$=\sum_{w\in\{0,1\}}Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},w}$
$=\sum_{w\in\{0,1\}}(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})(Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})$
$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
$=0(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})0=0$
where the penultimate equality follows from a combination of $($$\mathbf{QH_{1}}$)) and the fact that $(1,\dots,1)\notin R_{(1,0,\dots,0)}$.
This proves that every non-oracular quantum polymorphism of $\mathbb{O}_{(1,0,\dots,0)}$ preserves $R_{(1,0,0)}$ and thus
$\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})=\operatorname{\oplus Pol}(\mathbb{O}_{(1,0,0)}).$
It follows that every non-oracular quantum polymorphism of $\mathbb{O}_{(1,0,\dots,0)}$ is non-contextual, which completes the proof. ∎
For a generalisation of Lemma 51 to an arbitrary $t\in R_{1/k}$, we can pp-define $R_{t}$ in $\mathbb{O}_{(1,0,0\dots,0)}$ by
$R_{t}(x_{1},\dots,x_{k})\Leftrightarrow R_{(1,0,0\dots,0)}(x_{l},x_{1}\dots,x_{l-1},x_{l+1},\dots,x_{k})$
where $1\leq l\leq k$ is the singular non-zero component of $t$.
Corollary 52.
If $k\geq 3$ and $t\in R_{1/k}$, the non-oracular polymorphisms of $\mathbb{O}_{t}$ are non-contextual.
Now for the when $t\notin R_{1/k}$. We reduce it to the case from Corollary 52.
Lemma 53.
Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)\oplus_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
Proof.
Without loss of generality, assume that $t=t^{l}$ is the tuple with $l$ 1-entries at the start, one can do this from an arbitrary $t$ by just permuting the entries.
If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t\oplus_{2}t$ and $(1,\dots,1)=t\oplus_{2}\overline{t}$ are not in $R_{t}$. We first pp-define $\neq$ in $\mathbb{O}_{t}$ via
$x\neq y\Leftrightarrow R_{t}(x,\dots,x,y,\dots,y)$
where we fill the first $l+1$ entries with $x$’s. This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}^{n}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
$Q_{\mathbf{a},x}Q_{\mathbf{b},x}=Q_{\mathbf{a},x}\dots Q_{\mathbf{a},x}Q_{\mathbf{b},x}\dots Q_{\mathbf{b},x}=0.$
Let $t^{\prime}=(0,1\dots,1,0,\dots,0)$ ($t^{l}$ with the first entry switched). We define $R_{t^{\prime}}$ by the formula
$R_{t^{\prime}}(s_{1},\dots,s_{k})\Leftrightarrow\exists q\left(R_{t}(q,s_{2},\dots,s_{k})\wedge q\neq s_{1}\right).$
Assume we have tuples $\mathbf{s}_{1},\dots,\mathbf{s}_{k}\in\{0,1\}^{n}$ with $(\mathbf{s}_{1},\dots,\mathbf{s}_{k})$ in the $n$th direct power of $R_{t^{\prime}}$ and variables $x_{1},\dots,x_{k}\in\{0,1\}$ such that $(x_{1},\dots,x_{k})\notin R_{t^{\prime}}$. Then we have $\mathbf{q}\in\{0,1\}^{n}$ such that $(\mathbf{q},\mathbf{s}_{2},\dots,\mathbf{s}_{k})$ are elements of the $n$th power of $R_{t}$ and $\mathbf{q}\neq\mathbf{s}_{1}$ . Also, there is no $y\in\{0,1\}$ such that both $(y,x_{2},\dots,x_{k})\in R_{t}$ and $y\neq x_{1}$, i.e., we know that $(\neg x_{1},x_{2},\dots,x_{k})\notin R_{t}$.
Now let $Q\colon\mathbb{O}_{t}^{n}\to_{q}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
$Q_{\mathbf{s}_{1},x_{1}}\dots Q_{\mathbf{s}_{k},x_{k}}$ $=Q_{\mathbf{s}_{1},x_{1}}\left(\sum_{y\in\{0,1\}}Q_{\mathbf{q},y}\right)Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}}$
$=\sum_{y\in\{0,1\}}Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},y}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}}$
$=\sum_{y\in\{0,1\}}(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},y})(Q_{\mathbf{q},y}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})$
$=(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},x_{1}})(Q_{\mathbf{q},x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})+(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},\neg x_{1}})(Q_{\mathbf{q},\neg x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})$
$=(0)(Q_{\mathbf{q},x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})+(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},\neg x_{1}})(0)=0$
where the penultimate equality follows from $($$\mathbf{QH_{1}}$)). This proves that every non-oracular quantum polymorphism of $\mathbb{O}_{t^{l}}$ preserves $R_{t^{\prime}}$ and thus
$\operatorname{qPol^{no}}(\mathbb{O}_{t^{l}})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{t^{\prime}})=\operatorname{qPol^{no}}(\mathbb{O}_{t^{l-1}})$
by induction, we then get that
$\operatorname{qPol^{no}}(\mathbb{O}_{t})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})$
completing the proof after an application of Lemma 51.
Note that this theorem also covers the case when $t=(0,\dots,0)$ i.e. $R_{t}=R_{1/k}$ (or $l=0$). In this case $R_{t^{\prime}}$ as pp-defined above is itself $R_{(1,0,\dots,0)}$ and we are done immediately. ∎
Next, we argue that the only real cases we need to consider in the full theorem is the $\mathbb{O}_{t}$’s.
Lemma 54.
Every relational structure $\mathbb{A}$ with a relation $R$ that has any full binary projection has the property that $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Consider a relational structure with a relation $R$ that has a full binary projection on the components $\{i,j\}$. If this is the case, then in any power, any two elements appear together in some tuple of the direct power of RR.∎
Theorem 55.
For any relation $R\subseteq\{0,1\}^{k}$ the following statements are equivalent:
  1. (1)
    $R$ is not invariant under majority, majority preserves every proper projection of $R$, and every binary projection of $R$ is not full.
  2. (2)
    $R=R_{t}$ for some tuple $t$.
Proof.
2. implies 1. Note that translating a relation by a tuple does not change the size of the binary projections and does not change preservation by majority, since majority is a self-dual operation. It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in 1. are clear.
For 1. implies 2. suppose $R\subseteq\{0,1\}^{k}$ is not invariant under majority but majority does preserve every proper projection and no binary projection is full. Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority, we first show that $\{x\oplus_{2}t\mid x\in R_{1/k}\}\subseteq R$.
For $i\in[k]$, let $x\in R_{1/k}$ be the tuple with a 1 in the $i$th position. We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$. Now, since $t\notin R$; we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x\oplus_{2}t=r_{i}$, so $x\oplus_{2}t\in R$.
Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}=(1,\dots,1)\oplus_{2}t$ in exactly one argument, placing it in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$.
Assume that $r$ matches $\overline{t}$ in two arguments $1\leq i<j\leq k$. We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$ and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction. Additionally, $r$ cannot match $\overline{t}$ on zero arguments since then it would equal $t$, which is not in $R$. Thus every $r$ is in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$, completing the proof. ∎
We can now easily prove the general result.
Theorem 56.
If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Consider a relational structure $\mathbb{A}$; without loss of generality, every projection of a relation of $\mathbb{A}$ is also a relation of $\mathbb{A}$, since this does not change the hypothesis or the conclusion of the statement. Let $R$ be a relation of $\mathbb{A}$ that is not preserved by majority and of minimal arity. By Lemma 54, if a binary projection of $R$ is full then we are done, so assume that this is not the case. Then Theorem 55 implies that $R=R_{t}$ for some $k$-tuple $t$.
Finally, we know that $k\geq 3$ since every binary relation of arity 2 is preserved under majority and so, most of the result follows from Corollary 52 and Lemma 53. All that remains is for us to argue that Lemma 52 also covers the case where $\overline{t}\in R_{1/k}$.
This is since, $\mathbb{O}_{t}$ has ‘essentially’ the same polymorphisms as $\mathbb{O}_{\overline{t}}$ for any $t$. The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if and only if $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{\overline{t}}$. In particular $\mathbb{O}_{t}$ has contextual polymorphisms if and only if $\mathbb{O}_{\overline{t}}$ does. Thus we can assume that $\overline{t}\notin R_{1/k}$, completing the proof. ∎
8.2. The minimal contextual clone
In this section, we provide a clone with contextual polymorphisms that is contained within the quantum polymorphism clone of any Boolean relational structure with contextual polymorphisms.
In the previous section, we have seen that every boolean relational structure with contextual polymorphisms has non-full binary projections and contains majority in its (classical) polymorphism clone.
Let $\mathbb{B}$ be the structure $\mathbb{B}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$. In the classical setting, $\operatorname{Pol}(\mathbb{B})$ is the smallest clone that contains the majority operation. We first show that this property extends to oracular quantum polymorphisms.
Lemma 57.
Let $\mathbb{A}$ be a Boolean structure that admits majority as a polymorphism, and such that no binary projection of a relation of $\mathbb{A}$ is full. Then $\operatorname{qPol}(\mathbb{B})\subseteq\operatorname{qPol}(\mathbb{A})$.
Proof.
Consider a relational structure $\mathbb{A}:=(\{0,1\};R)$ that is preserved by majority, then it is well known that $R$ (of arity $r$) can be pp-defined in $\mathbb{A}_{2}:=(\{0,1\};\{\pi_{i,j}(R)\mid 1\leq i,j\leq r\})$ via the conjunction of its binary projections $\bigwedge_{1\leq i,j\leq r}\pi_{\{i,j\}}(R)(x_{i},x_{j})$. We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to\mathbb{A}_{2}$. Assume that $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})$ is in the $n$th direct power of $R$ while in $(x_{1},\dots,x_{r})$, we have indices $i,j$ such that $(x_{i},x_{j})\notin\pi_{\{i,j\}}(R)$. Then
$Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=(Q_{\mathbf{a}_{i},x_{i}}Q_{\mathbf{a}_{j},x_{j}})Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=(0)Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=0$
where the second equality is by duplicating the $Q_{\mathbf{a}_{i},x_{i}}$ and $Q_{\mathbf{a}_{j},x_{j}}$ projectors and then using property $($$\mathbf{QH_{2}}$)) to move them to the front. Moreover, if $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)). $Q$ is then an oracular quantum polymorphism of $\mathbb{A}$ and so $\operatorname{qPol}(\mathbb{A}_{2})\subseteq\operatorname{qPol}(\mathbb{A})$.
By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty!) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$. We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\mathbb{B}$.
Let $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$. We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $($$\mathbf{QH_{1}}$)) holds.
Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute, so that $($$\mathbf{QH_{2}}$)) holds. ∎
Now we show that this clone actually contains some contextual polymorphisms.
Theorem 58.
For $n\leq 3$, every quantum polymorphism $Q\colon\mathbb{B}^{n}\to\mathbb{B}$ is non-contextual. For all $n\geq 4$, $\mathbb{B}$ has contextual quantum polymorphisms of arity $n$.
Proof.
We start by proving that every quantum polymorphism $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ is non-contextual for $n\leq 3$. For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\mathbb{B}^{n}$. If $(S,T)$ does not belong to the relation $S_{00}$ in $\mathbb{B}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$. Similarly, for it to not belong to the relation $S_{11}$ in $\mathbb{B}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$. If $n\leq 3$, there remains at one most other element $k$ in $\{1,\dots,n\}\setminus\{i,j\}$. If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\mathbb{B}^{n}$, so we are done.
For the $n\geq 4$ case we define a concrete quantum polymorphism for $n=4$. We provide Figure 3 as a visual aid.
0000100001000010000111001010100101100101001111101101101101111111
Figure 3. Simplified illustration of $\mathbb{B}^{4}$. Black lines indicate the poset $S_{10}$. The blue (resp. red) dotted lines show parts of $S_{00}$ (resp. $S_{11}$); this relation additionally relates $(0,0,0,0)$ (resp. $(1,1,1,1)$) to everything and relates two nodes when they are a bit-flips of each other (see coloring).
We define $Q\colon\mathbb{B}^{4}\to_{q}\mathbb{B}$ by listing the projectors. Let $T\subseteq[4]$. If $|T|\geq 3$, define $Q_{T}=\operatorname{id}$. Pick now arbitrary matrices $A,B,C$. Define $Q_{\{1,2\}}:=A$, $Q_{\{1,3\}}=B$, and $Q_{\{1,4\}}=C$. In all the other cases, define $Q_{T}=\operatorname{id}-Q_{[4]\setminus T}$. Note that if at least two of the matrices $A,B,C$ do not commute, then $Q$ is contextual.
It remains to prove that $Q$ preserves the relations of $\mathbb{B}^{4}$. Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and $x,y\in\{0,1\}$ with $(x,y)\notin S_{11}$. If $S\neq\overline{T}$, we are in one of 2 cases; either $S$ or $T$ is the empty set or $|S|=|T|=1$. In both cases at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately. If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\overline{S}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done. The case of $S_{00}$ is symmetric.
Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$. If both $|S|$ and $|T|$ do not equal $2$ we know $Q_{S},Q_{T},(\operatorname{id}-Q_{S}),(\operatorname{id}-Q_{T})\in\{0,\operatorname{id}\}$ and we immediately get that $Q_{S},{\operatorname{id}-Q_{T}}]=0$. The final case is when one of $|S|$ or $|T|$ equal $2$, here we get that either
$Q_{S}(\operatorname{id}-Q_{T})=0(\operatorname{id}-Q_{T})=0$
or,
$Q_{S}(\operatorname{id}-Q_{T})=Q_{S}(\operatorname{id}-\operatorname{id})=0$
showing preservation of the relations. ∎
9. Acknowledgments
The first and last author are thankful for the attentive listening and comments by Peter Zeman, to whom they explained the definition of quantum polymorphisms and the main results in this paper concerning the characterisation of the existence of commutativity gadgets on the occasion of the CSP World Congress in September 2025. The second and last author acknowledge funding from the project “Hamburg Quantum Computing”, co-financed by ERDF of the European Union and by the Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB).